### v1

# Connecting the Dots: A Chain-of-Collaboration Prompting Framework for LLM Agents

Abstract
Large Language Models (LLMs) have demonstrated impressive performance in executing complex reasoning tasks. Chain-of-thought effectively enhances reasoning capabilities by unlocking the potential of large models, while multi-agent systems provide more comprehensive solutions by integrating collective intelligence of multiple agents. However, both approaches face significant limitations. Single-agent with chain-of-thought, due to the inherent complexity of designing cross-domain prompts, faces collaboration challenges. Meanwhile, multi-agent systems consume substantial tokens and inevitably dilute the primary problem, which is particularly problematic in business workflow tasks. To address these challenges, we propose Cochain, a collaboration prompting framework that effectively solves business workflow collaboration problem by combining knowledge and prompts at a reduced cost. Specifically, we construct an integrated knowledge graph that incorporates knowledge from multiple stages. Furthermore, by maintaining and retrieving a prompts tree, we can obtain prompt information relevant to other stages of the business workflow. We perform extensive evaluations of Cochain across multiple datasets, demonstrating that Cochain outperforms all baselines in both prompt engineering and multi-agent LLMs. Additionally, expert evaluation results indicate that the use of a small model in combination with Cochain outperforms GPT-4.
**footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Instruction
Large Language Models (LLMs) have demonstrated impressive performance in language understanding and complex reasoning tasks [51, 5, 21, 8], with prompt engineering playing a crucial role in enhancing their reasoning capabilities. Prompt engineering represented by Chain-of-Thought prompt LLMs in ways that facilitate coherent and step-by-step reasoning processes, endowing agents with depth of thought [7, 57, 66]. To enhance agent collaborative capabilities, various multi-agent collaborative frameworks have been developed, providing agents with breadth of thought. These frameworks encode carefully designed agent profiles and collaborative mechanisms into prompts, yielding favorable outcomes across domains such as healthcare [50, 29, 38], education [17, 44, 22], law [16], and finance [64]. Due to their exceptional problem-solving and collaborative abilities, multi-agent systems are considered a promising pathway toward Artificial General Intelligence (AGI) [9, 81]. However, this raises a critical question: Is a multi-agent LLM system all you need?
Evidently not, as both single-agent and multi-agent systems exhibit limitations [53]. As illustrated in Figure 1, the inherent complexity in designing cross-domain prompts [47] means that single agents employing prompt engineering lack constraint awareness and face challenges in cross-domain collaboration (under-collaboration). Simultaneously, multi-agent systems face challenges from expensive token consumption and extensive inference time [20, 60], with inter-agent communications often ineffectively utilized [72]. More importantly, existing research primarily focuses on maximizing collaboration, extensively exploring how multi-agent systems can improve decision quality [73, 43], make safe decisions [33, 42], and solve problems [78, 34], while paying insufficient attention to the negative effects of excessive agent collaboration. When all agents simultaneously participate in every decision-making process, core issues are frequently overwhelmed by non-critical information, leading not only to a decline in answer quality but also to a preference for consensus-driven answers over accurate ones [20], particularly in business workflow tasks. To our best knowledge, no existing literature has systematically analyzed and defined this “over-collaboration” phenomenon or proposed targeted solutions. This prompts us to consider: Why not combine the reasoning capabilities of chain-of-thought prompting with the collaborative nature of multi-agent systems?
Inspired by these challenges, we propose a collaborative prompting framework, Cochain, designed to enhance LLM agents’ collaborative capabilities in business workflow tasks at low cost, while mitigating over-collaboration through indirect rather than direct agent participation in decision-making. As shown in Figure 2, Cochain comprises two primary components: collaborative knowledge graph and prompts tree. We mine agent knowledge through counterfactual reasoning and refine it into a knowledge graph integrated with explicit dataset knowledge. To mitigate the impact of irrelevant knowledge on model performance [79], we introduce causal chain structure to improve cross-stage reasoning capabilities. Inspired by how the human brain integrates fragmented knowledge into coherent thought [15], we propose prompts tree that distills prompts from agent responses. By retrieving from the prompts tree, agents can automatically acquire chain-like cross-stage prompts, avoiding the high cost of manual prompt design while effectively utilizing inter-agent communications.
Figure 1: The impact of agent collaboration degree on problem-solving effectiveness. Under-collaboration ignores constraints from other stages, while over-collaboration dilutes main issues with irrelevant information (underlined in our examples). Agents using Cochain achieve balance by considering constraints while maintaining focus on key problems, effectively connecting the dots.
The main contributions are summarized as follows:
  • $\bullet$
    Effective Collaboration: We propose an agent collaborative prompting framework, Cochain, that equips agents with both depth and breadth of thinking, effectively alleviating the problems of over-collaboration and under-collaboration.
  • $\bullet$
    Cost-Effective yet Efficient: Our extensive experimental results demonstrate that Cochain exhibits significant improvements over all baselines across different model backbones and three business workflow datasets. Cochain possesses reasoning speeds comparable to single-agent systems and even lower inference costs, and surpasses existing multi-agent systems in capability, achieving an optimal balance between performance and efficiency.
  • $\bullet$
    The Small + Cochain Outperforms the Large: Evaluations by domain experts have further confirmed that the combination of the small model and our framework outperforms the large model. To support further research on over-collaboration and business workflow tasks, we will publicly release our collaborative knowledge graph and high-quality datasets.
2 Related Work
Figure 2: Cochain connects business workflow stages. In part (a), agents address counterfactual inquiries, with answers undergoing interactive iteration. In part (b), a prompts tree is built. Part (d) presents our framework: an agent uses the prompts tree and a collaborative knowledge graph built from counterfactual answers and original data, achieving more collaborative results than part (c).
2.1 Knowledge Graph Augmented LLMs
LLM reasoning capabilities are critical for high-quality responses [28, 49, 30]. Knowledge graphs, with their structured, explicit, and interpretable nature [75, 40, 80], offer clear representations and transparent reasoning paths, effectively guiding LLMs in deeper reasoning and mitigating limitations [36, 67, 56]. For domain-specific LLMs, knowledge graphs enable continuous updating of knowledge frameworks to accommodate evolving information [27, 37, 31]. Currently, MindMap [59] applies knowledge graphs in medicine, while ChatLAW [16] employs them in legal domains for process simulation. Business workflows, due to their multi-stage nature and cross-domain integration requirements, face the unique challenge of multi-stage knowledge fusion. To prevent performance degradation from irrelevant knowledge [39], we propose a causal chain optimizing knowledge selection, enhancing understanding of cross-stage relationships while filtering out noise.
2.2 Prompt Engineering
Prompt engineering aims to guide LLMs to maximize their potential through carefully designed prompts [11, 47, 76]. Its core enables models to adapt to new tasks by leveraging contextual learning and following instructions [10, 32]. Research shows well-designed prompts significantly enhance model performance across tasks, particularly complex ones [57, 55, 77, 65]. CoT [57] enables models to generate intermediate reasoning steps through simple prompts, improving performance. ToT [66] explores coherent text units as intermediate problem-solving steps. CPO [71] leverages non-optimal reasoning paths generated through tree search, speeding up inference. However, existing prompt engineering methods lack sustained constraint awareness and neglect cross-domain collaboration. Cochain enables LLM agents to address this challenge while balancing interdependent constraints and implementation feasibility.
2.3 Collaboration of LLM Agents
In research based on LLMs, many studies focus on solving problems through the collaboration of multiple agents [4, 20, 23, 74, 24]. These approaches adopt diverse collaboration paradigms. Some leverage discussion  [14, 50, 46] and debate [20, 63, 12] frameworks to enhance reasoning capabilities. Others implement hierarchical structures with specialized roles [68]. Several develop sequential pipelines or tree-structured architectures [72, 74]. These collaborative frameworks have demonstrated effectiveness across various domains, including medicine [50], long-text processing [72, 74, 48, 13], social simulation [69, 58], and code intelligence [54, 26]. Different from the above works, Cochain relies on knowledge fusion rather than token-intensive role-playing mechanisms. To the best of our knowledge, we are the first to address LLM agent collaboration for business workflow tasks, offering an efficient solution that significantly reduces computational overhead while maintaining strong performance in real-world scenarios.
3 Methods
3.1 Collaborative Knowledge Graph
The construction of the collaborative knowledge graph relies on a large amount of stage-specific knowledge. We adopt the Self-Instruct [55] to generate vertical domain instruction datasets $D_{i}=\{(X_{ij},Y_{ij})\}_{j=1}^{n_{i}}$ for fine-tuning LLMs, where $X_{ij}\in X_{i}$ is the input and $Y_{ij}\in Y_{i}$ is the corresponding output. The datasets are subject to data cleaning and triplet extraction to acquire knowledge for constructing the explicit knowledge graph $\mathcal{KG}_{\text{explicit}}$:
$\mathcal{KG}_{\text{explicit}}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{n_{i}}\text{ExtractTriples}(X_{ij},Y_{ij})$ (1)
In order to leverage the capabilities of other business workflow agents, we need to identify the tacit variables that play a critical role in the agent’s responses. Since extracting the internal tacit variables of agents is challenging, we focus on the extraction of tacit knowledge from the hidden layers. To achieve this, we introduce counterfactual reasoning. Specifically, we perform causal reasoning, adversarial reasoning, substitution reasoning, extreme counterfactual reasoning, and backward causal reasoning on the questions within the datasets. For each sample $(X_{ij},Y_{ij})$, we generate the corresponding counterfactual input $\tilde{X}_{ij}$:
$\tilde{X}_{ij}=\text{GenerateCounterfactual}(X_{ij})$ (2)
We then input these questions into the relevant vertical domain agents to generate answers. To represent the output of tacit knowledge influenced by tacit variables, we introduce the tacit variable $\theta_{ij}$, which represents the tacit knowledge or state inside the model and plays a key role in generating counterfactual outputs. The tacit knowledge $h_{ij}$ is the model’s hidden state under the influence of the tacit variable $\theta_{ij}$, which acts as an intermediary in the process of generating counterfactual outputs. The tacit knowledge $h_{ij}$ is influenced by the tacit variable $\theta_{ij}$ where $h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$. The tacit knowledge $h_{ij}$ further influences the generation of the counterfactual output $\tilde{Y}_{ij}$, where $\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The final generation process is expressed as (Appendix A):
[…]
In this process, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ represents the probability of generating the counterfactual output given the tacit knowledge $h_{ij}$ and the counterfactual input $\tilde{X}_{ij}$. $P(h_{ij}\mid\theta_{ij})$ represents the generative distribution of the tacit knowledge $h_{ij}$ given the tacit variable $\theta_{ij}$, and $P(\theta_{ij}\mid\tilde{X}_{ij})$ represents the prior distribution of the tacit variable $\theta_{ij}$ given the counterfactual input $\tilde{X}_{ij}$. Subsequently, these answers are evaluated by a general-purpose LLM, and the evaluation results are classified into three levels: reasonable, ambiguous, and unreasonable. The general-purpose model provides feedback on evaluation opinions and suggestions, which are then used by the vertical domain agent to generate another response. This process iterates until a reasonable answer is obtained.
During knowledge distillation, we apply causal enhancement, with special attention to causal keywords such as “depends on”, “relies on”, and “applies to”. Using the estimated tacit variable $\theta_{ij}$, tacit knowledge is extracted from the counterfactual input-output pairs $(\tilde{X}_{ij},\tilde{Y}_{ij})$ completed through iteration, and a tacit knowledge graph $\mathcal{KG}_{\text{tacit}}$ is constructed:
$\mathcal{KG}_{\text{tacit}}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{n_{i}}\text{ExtractTriples}(\tilde{X}_{ij},\tilde{Y}_{ij},\theta_{ij})$ (4)
Finally, by integrating the explicit knowledge graph and the tacit knowledge graph, we construct a collaborative knowledge graph $\mathcal{KG}$ that contains both stage-specific knowledge and inter-stage connectivity:
$\mathcal{KG}=\mathcal{KG}_{\text{explicit}}\cup\mathcal{KG}_{\text{tacit}}$ (5)
3.2 Causal Chain
To enable the agent to organize the causal relationships between various knowledge stages and generate more accurate and coherent responses, rather than simply listing knowledge, we introduce a causal chain mechanism. When a user submits a request to the agent, the request is first decomposed, and keywords are extracted. These keywords are then input into the knowledge graph to match relevant nodes.
Figure 3: Retrieving relevant knowledge from the collaborative knowledge graph and the process of constructing causal chains.
Specifically, as shown in Figure 3, we use a pre-trained text encoder [45] to convert the extracted keywords and their related nodes into vector representations. For the vector representation $v_{k}$ of keyword $k$ and the vector representation $v_{n}$ of node $n$ in the knowledge graph, we compute the cosine similarity between these vectors to identify the most relevant stage knowledge:
$\text{Sim}(v_{k},v_{n})=\frac{v_{k}\cdot v_{n}}{\|v_{k}\|\|v_{n}\|}$ (6)
By setting a similarity threshold $\delta$, we select knowledge nodes that are highly related to the semantics of the keywords. For these nodes, we extend their one-hop neighbors to introduce additional associated knowledge and construct a triple-based knowledge structure to enhance the agent’s understanding capability. To address the challenge of cross-stage integration, we adopt a multi-hop neighbor exploration method to build causal chains. By leveraging stage-specific knowledge as a bridge, this approach effectively explains inter-stage knowledge relationships. Specifically, within the stage set $X=\{x_{1},x_{2},x_{3},\dots\}$, for a node $A$ with high semantic similarity, where $A\in x_{1}$, we expand its one-hop neighbor node $B$, where $B\in x_{1}\cap x_{i}$ ($i\neq 1$), and use this as a bridge to explore nodes $C$ from other stages, where $C\in x_{i}$ ($i\neq 1$), to construct the causal chain:
[…]
Using the causal chain mechanism, the agent is able to comprehend and identify the causal relationships between the knowledge we provide. This enables the agent to consider the interdependencies between different components when generating responses to user queries.
Table 1: Performance comparison of different model backbones and multi-agent baselines across three business domains. The performance of our proposed method and all baselines. BERTScore F1 is abbreviated as BS-F. We highlight the best and second-best results.
Backbone Baseline Automotive Pharmaceutical E-commerce
BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L
Qwen2-7B PMC [68] 65.81 13.25 18.37 68.55 17.34 25.93 70.88 24.01 29.91
MedAgents [50] 64.94 12.47 18.34 65.96 10.63 21.13 69.57 17.48 33.62
Debate(short) [20] 65.32 12.66 17.81 70.57 15.78 23.03 72.93 20.85 28.23
Debate(long) [20] 65.56 12.92 18.03 61.14 9.58 12.23 72.89 20.81 28.04
CoA [72] 71.72 22.31 24.01 77.26 29.92 33.70 79.20 38.08 39.02
Cochain 75.39 28.17 28.87 78.48 33.98 35.03 79.91 37.08 40.06
DeepSeek-R1-7B PMC [68] 65.99 11.09 16.00 67.95 23.02 28.18 67.52 16.52 23.53
MedAgents [50] 64.94 12.47 18.34 70.44 24.77 30.73 67.19 14.60 21.18
Debate(short) [20] 66.20 11.45 17.39 71.52 27.03 30.56 70.72 22.92 26.99
Debate(long) [20] 65.94 11.45 17.21 71.77 27.19 30.39 70.68 22.93 27.16
CoA [72] 71.41 12.12 17.55 77.29 38.14 35.43 80.33 43.81 41.54
Cochain 73.19 21.73 24.01 78.97 37.94 36.75 82.68 50.30 47.24
Claude-3.5-haiku PMC [68] 64.11 10.86 10.71 66.51 12.33 13.56 66.96 17.17 16.28
MedAgents [50] 64.45 11.12 11.09 66.23 14.02 14.32 68.11 18.18 16.76
Debate(short) [20] 64.67 11.27 13.44 69.16 14.81 16.06 68.62 19.77 18.25
Debate(long) [20] 65.24 10.88 12.76 67.64 15.32 15.44 68.79 18.06 20.08
CoA [72] 63.39 8.99 10.21 65.22 13.66 14.27 66.92 15.20 15.01
Cochain 69.61 18.73 17.30 72.87 23.59 19.67 71.36 20.15 20.78
3.3 Prompts Tree
In business workflow tasks, the interconnections between various stages are crucial for effective collaboration. To achieve this, we propose an innovative prompt distillation method aimed at constructing a prompts tree. This method derives stage-specific Q&A from the seed Q&A using a fragmentation approach, where each node at every level of the tree corresponds to prompts from different stages. By querying the prompts tree, it is possible to obtain prompt information that spans the entire business workflow. Specifically, we start with seed Q&A in Stage 1 and utilize the agent from Stage 1 to distill solution-oriented prompts from the answers. The most effective $m$ prompts are selected through self-evaluation. These $m$ prompts are then used to generate new question templates, focusing primarily on how to implement or solve the prompts distilled from the previous stage in the next stage. Each question is answered, forming a new seed Q&A. These Q&A are further involved in the prompt distillation of the current stage, driving the content to extend to the next stage.
In this process, the user need is treated as the root node, with prompts in Stage 1 derived as child nodes based on the user need. These Stage 1 prompts further generate Q&A in Stage 2, from which Stage 2-specific prompts are distilled as child nodes of the Stage 1 prompts. In this way, we continuously build and maintain a prompts tree. By retrieving the prompts tree, we automatically generate a prompt chain that covers the entire business workflow, thus avoiding the difficulties associated with manually constructing prompts. These prompts will guide large language models to generate responses that comprehensively consider the content of each stage, ensuring high collaboration across the stages.
4 Experiments
In our experiments, we evaluate our collaboration framework and answer the following questions: (1) To what extent does Cochain enhance inter-agent collaboration capabilities? (2) How does Cochain affect agent performance across specialized and cross-domain collaboration tasks? (3) Which key design elements of Cochain are most critical for enhancing collaboration capabilities?
4.1 Experiments Setup
Datasets.
To comprehensively evaluate our method’s impact on business workflow collaboration, we construct novel datasets. Given the scarcity of existing collaborative datasets and their limited ability to reflect complex cooperation patterns in authentic business environments, we develop three business workflow datasets of increasing complexity: e-commerce, pharmaceutical, and automotive workflows. Detailed datasets construction methodology is provided in Appendix B. Notably, the automotive workflow test set underwent rigorous review and refinement by domain experts from a major automotive manufacturer, ensuring its accuracy and practical relevance.
Metrics.
BERTScore [70] utilizes contextual word embeddings to evaluate semantic similarity, while GLEU [62] emphasizes grammatical and semantic quality, and ROUGE-L [35] assesses deep semantic alignment, particularly for extended texts. We employ these metrics as primary indicators of semantic similarity between generated and reference answers. For comprehensive experimental results across additional metrics, refer to Appendix C.
LLMs.
We utilize fifteen LLMs as the backbone of Cochain. With both proprietary and open-source models, we examine the differences between reasoning-capable models (DeepSeek-R1-7B [18]) and standard models (Qwen2-7B [2]). We also investigate using identical versus mixed models (Llama2-7B [52], GLM4-9B, Qwen2-7B, DeepSeek-7B) across workflow stages. All open-source models undergo domain-specific fine-tuning to serve as specialized agents in target domains. For proprietary models, we assess GPT-3.5-turbo, GPT-4o [3], and Claude-3.5-haiku [1] performance.
Baselines.
Our baseline selection encompasses the predominant multi-agent collaboration methodologies (Appendix B). Specifically, PMC [68] implements multi-agent collaboration through Hierarchical Planning. MedAgents [50] facilitates collaborative decision-making via discussion and voting mechanisms. Debate [20] enables collaboration through argumentative discourse, modulating agent confidence through varying debate durations. CoA [72] establishes chain-based collaboration where agents interact through dedicated CU. Besides, we compare the performance of single-agent prompt engineering techniques, namely CoT [57] and ToT [66].
Figure 4: (a) Expert evaluation & (b) GPT-4 scores.
Figure 5: Reasoning time differential.
Table 2: The performance of GPT-3.5-turbo and GPT-4o when using CoT, ToT, and Cochain.
Backbone Method BS-F GLEU ROUGE-L
GPT-3.5-turbo + IO 72.78 15.28 22.06
+ COT [57] 71.28 10.91 17.73
+ TOT [66] 66.79 10.15 17.92
+ Cochain 76.92 28.68 30.73
GPT-4o + IO 71.69 22.71 23.48
+ COT [57] 72.70 24.03 24.44
+ TOT [66] 66.29 12.54 17.82
+ Cochain 74.51 25.17 25.22
Table 3: Result of different skip stages on evaluation metrics.
Skip Stage BS-F GLEU ROUGE-L
zero / 75.39 28.17 28.87
one $S_{1}\rightarrow S_{3}\rightarrow S_{5}$ 74.00 24.66 25.18
$S_{2}\rightarrow S_{4}$ 74.52 25.57 26.18
$S_{3}\rightarrow S_{5}$ 74.18 25.45 25.95
two $S_{1}\rightarrow S_{3}$ 74.27 25.26 25.74
$S_{2}\rightarrow S_{4}$ 74.07 25.00 25.82
4.2 Enhancing Collaborative Reasoning Capabilities in LLM Agents with Cochain
Overall Results of Cochain.
Table 1 demonstrates Cochain’s performance on business workflow tasks across three diverse language models. Results show Cochain consistently outperforms all baselines by significant margins across all datasets. Cochain also exhibits substantial advantages in computational efficiency, with shorter reasoning times, fewer output tokens, and lower operational costs compared to other baselines (Appendix C.1). Figure 7 reports consistent performance gains achieved by Cochain as model parameters increase, demonstrating its strong scalability potential with more powerful foundation models. As shown in Figure 5, the reasoning time differential between thinking and non-thinking models across increasing reasoning rounds reveals that with PMC, this differential grows dramatically as rounds increase, while with Cochain, it remains relatively moderate. This demonstrates Cochain’s superior efficiency in handling complex reasoning tasks, with its performance advantage becoming increasingly pronounced as reasoning complexity increases.
Comparative Analysis of Cochain Against Prompting Engineering Approaches.
Table 3 presents a comparative evaluation of Cochain against established prompting methodologies. The results show that while CoT and ToT approaches enhance reasoning capabilities, they exhibit limited collaborative advantages compared to our framework. Specifically, these traditional techniques focus on isolated reasoning paths, whereas Cochain enables agents to incorporate both upstream and downstream factors into decision-making, yielding more comprehensive solutions for complex, interdependent problems across multiple domains of expertise.
Qualitative Evaluation.
Given the absence of established metrics for evaluating LLM agent collaboration, we conduct a two-fold validation approach. First, we recruit five experts from a major automotive manufacturer for evaluation. These experts judge Qwen2-7B with Cochain to outperform GPT-4 in approximately 60% of cases (Figure 5(a)). We also employ GPT-4 for quantitative assessment across five scoring rounds, comparing performance with and without fine-tuning and our framework. Figure 5(b) shows models enhanced with Cochain (Raw* and Finetune*) demonstrate significantly higher median scores, with the fine-tuned version achieving nearly optimal performance (median >9.7). These evaluations provide compelling evidence for Cochain’s real-world effectiveness.
Figure 6: Parameter analysis results. In the first five plots, the top-n value is fixed, while in the last plot, the number of prompts is fixed.
Figure 7: Comparative analysis of metrics scaling with increasing model size from DeepSeek-R1 7B to 70B.
Table 4: Results of multi-stage collaboration performance on Automotive datasets, highlighting fine-tuned results, Cochain application, and improvements()($\uparrow$)( ↑ ). The experimental results reflect two different scenarios in the workflow. The score is derived from GPT-4’s rating of the final solution (10-point scale), which is synthesized by integrating the answers from each stage.
Stage Method Backbone BS-F GLEU ROUGE-L Backbone BS-F GLEU ROUGE-L
$S_{1}$ Finetuned Qwen2-7B 71.60 22.59 24.51 Llama2-7B 70.68 18.79 21.49
+ Cochain 74.65 26.28 25.97 71.46 20.96 21.58
$\uparrow$ +3.05 +3.69 +1.46 +0.78 +2.17 +0.09
$S_{2}$ Finetuned 70.44 19.37 21.70 GLM4-9B 69.34 16.30 19.76
+ Cochain 74.65 26.57 25.74 73.19 22.33 24.06
$\uparrow$ +4.21 +7.20 +4.04 +3.85 +6.03 +4.30
$S_{3}$ Finetuned 70.23 19.39 22.39 Qwen2-7B 70.28 18.95 22.38
+ Cochain 74.84 27.39 26.29 74.51 25.76 26.00
$\uparrow$ +4.61 +8.00 +3.90 +4.23 +6.81 +3.62
$S_{4}$ Finetuned 69.42 18.71 21.90 DeepSeek-7B 68.96 13.06 18.70
+ Cochain 75.39 29.28 27.94 71.03 14.39 18.89
$\uparrow$ +5.97 +10.57 +6.04 +2.07 +1.33 +0.19
Score 6.3 9.2 +2.9 6.5 9.6 +3.1
4.3 Cochain Efficacy Across Specialized and Cross-Domain Tasks
To investigate Cochain’s impact on specialized and cross-domain tasks, we conducted experiments on automotive industry data using two approaches: a multi-domain model (single Qwen2-7B fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), with testing conducted on 4 sub-datasets. Table 4 shows that Cochain significantly improves both approaches, with particularly strong GLEU score gains. The specialized configuration demonstrates Cochain’s ability to orchestrate effective collaboration among domain-expert agents. GPT-4 evaluation scores, assessing comprehensive solutions resulting from four stages, confirm substantial quality improvements for both approaches. Additional experiments with specialized models of the same architecture fine-tuned for different domains are presented in Appendix C.3, further validating Cochain’s effectiveness in coordinating expert agents toward coherent solutions.
Table 5: Result of ablation study. $\mathcal{KG}$ represents the knowledge graph. We highlight the most and second-most efficient modules.
Backbone Method Automotive Pharmaceutical E-commerce
BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L
Qwen2-7B Cochain 75.39 28.17 28.87 78.48 33.98 35.03 79.91 37.08 40.06
w/o $\mathcal{KG}$ 75.21 27.81 28.74 76.34 25.92 28.09 76.72 32.40 32.74
w/o Causal Chain 75.20 27.55 28.42 76.20 25.38 27.76 76.86 33.30 32.93
w/o Prompts Tree 73.08 22.45 24.56 76.19 26.84 28.91 76.59 31.83 32.38
DeepSeek-R1-7B Cochain 73.19 21.73 24.01 78.97 37.94 36.75 82.68 50.30 47.24
w/o $\mathcal{KG}$ 72.41 19.71 22.63 78.22 37.33 36.40 77.36 36.54 35.85
w/o Causal Chain 72.67 19.96 22.86 78.41 37.36 36.74 77.60 36.30 35.74
w/o Prompts Tree 72.27 16.54 20.60 76.96 37.98 35.61 79.02 41.03 38.84
Claude-3.5-haiku Cochain 69.61 18.73 17.30 72.87 23.59 19.67 71.36 20.15 20.78
w/o $\mathcal{KG}$ 68.98 16.63 16.31 70.93 16.42 18.45 71.13 18.85 20.22
w/o Causal Chain 69.33 17.31 16.65 71.33 17.43 18.78 71.23 18.89 20.45
w/o Prompts Tree 65.79 12.84 13.71 70.85 16.24 18.31 69.76 15.55 18.59
4.4 Key Design Factors and Hyperparameter Influence in Cochain
Ablation Study.
To validate Cochain’s effectiveness, we conduct ablation experiments with three variants: (1) w/o Knowledge Graph; (2) w/o Causal Chain; and (3) w/o Prompts Tree. As shown in Table 5, all variants exhibit performance degradation across all metrics, confirming each component’s importance. Notably, removing the prompts tree causes the most significant drop, highlighting its critical role in collaboration(Appendix C.4). The knowledge graph and causal chain components also prove essential, demonstrating their complementary functions in the reasoning process.
Hyper-parameter Sensitivity Analysis.
We investigated key hyperparameters by fixing the number of nodes ($n$) from the knowledge graph while varying the prompt count. Figure 7 shows that for any fixed $n$, performance follows an inverted U-shaped curve, indicating an optimal range rather than a monotonic relationship. When fixing prompt count and varying $n$, no consistent pattern emerged, suggesting prompt design exerts greater influence than knowledge quantity, aligning with our ablation results. Table 3 demonstrates that skipping any stage ($S_{k}$) in the prompts tree consistently reduces performance, confirming each stage’s indispensable role in the collaborative reasoning process.
4.5 Overcollaboration is More Serious than Undercollaboration
Business workflow tasks are particularly susceptible to the “overcollaboration” phenomenon. When multiple agents collaborate, this manifests as responses deviating from core issues and reducing answer quality. Comparative analysis of Table 1, Table 4, and additional single-agent experimental results (Appendix C.2) demonstrates that, despite consuming significantly more computational resources, overcollaboration performs substantially worse than undercollaboration. Overcollaboration’s challenging nature is its covertness—systems maintain high activity and apparent collaboration, making the problem hard to detect and rectify promptly. While task decomposition, such as approaches like PMC, and summarization, such as strategies like CoA, are considered effective methods for focusing on core issues [61, 14], Cochain’s collaboration mechanism exhibits superior performance in controlling excessive collaboration, enabling a more precise focus on critical task requirements.
5 Conclusion
In this paper, we propose Cochain, a chain-of-collaboration prompting framework designed to enhance the collaboration of LLM agents in workflow tasks. Cochain combines the advantages of both single-agent and multi-agent systems, achieving depth in thought and breadth in collaboration. It is characterized by low operational cost, high inference speed, strong interpretability, and robust reasoning chain resilience. Cochain introduces a prompts tree structure for reasoning, enabling collaborative inference paths among multiple agents on-chain. By constructing a cross-node tacit knowledge graph, it improves relational reasoning and reduces hallucinations during collaboration. Additionally, the causal chain mechanism enhances interpretability. Extensive experiments demonstrate that Cochain yields effective improvements in scenarios suffering from both undercollaboration and overcollaboration, excelling in specialized and cross-domain tasks. Expert evaluations also indicate that combining smaller models with Cochain surpasses GPT-4 in workflow tasks.
Appendix A Proof: Tacit Knowledge & Counterfactual Output Model
Given the inherent difficulty in directly extracting hidden layer knowledge from agents operating within complex systems, our approach introduces counterfactual reasoning as a means to probe and subsequently infer tacit knowledge. This section elucidates the probabilistic formulations developed to model these inferred internal aspects of each agent. It details the conceptualization of tacit knowledge, the generation of counterfactual outputs by the agent, and the overall generative process used to understand these unobservable agent workings.
The foundation of our approach to modeling tacit knowledge within each agent lies in its conditional generation based on latent tacit variables. We define $h_{ij}$ as the tacit knowledge specific to a sample $(X_{ij},Y_{ij})$ processed by a particular agent. Within an agent, $h_{ij}$ represents an internal state or representation that is not directly accessible. While $h_{ij}$ is an abstract construct for modeling these unobservable aspects, for illustrative purposes, one might conceptualize it as akin to intermediate computational states, learned feature representations, or internal configurations within the Agent that influence its output generation. Concurrently, $\theta_{ij}$ is defined as a latent tacit variable associated with the same sample and agent. This variable $\theta_{ij}$ is also an abstract construct, encapsulating underlying factors hypothesized to influence $h_{ij}$, such as specific input characteristics processed by the Agent, its operational modes, or conditioning contexts it operates under. The introduction of $\theta_{ij}$ and $h_{ij}$ is an attempt to create a probabilistic model for the Agent’s internal, unobservable precursors to decision-making or generation.
A core assumption of our model is that this tacit knowledge $h_{ij}$ is not deterministically derived but is rather a stochastic realization conditioned on $\theta_{ij}$. This is particularly relevant for complex Agents, where output generation can be influenced by internal stochastic mechanisms or intricate decision processes. This probabilistic relationship is formally expressed as:
$h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$ (8)
This notation signifies that $h_{ij}$ is a random variable whose probability distribution $P(h_{ij}\mid\theta_{ij})$ is conditional upon the specific value or state of $\theta_{ij}$. The nature of this distribution would be specific to the agent and reflects the inherent complexities and uncertainties in how its internal states are formed. For instance, $\theta_{ij}$ might represent an abstract control signal or a high-level interpretation of input patterns by the agent, and $P(h_{ij}\mid\theta_{ij})$ would describe the distribution of possible internal representations resulting from it. This formulation acknowledges that even with a defined influencing factor $\theta_{ij}$, the precise tacit knowledge $h_{ij}$ (the agent’s internal processing focus or state) can vary. The characteristics of $P(h_{ij}\mid\theta_{ij})$ can be explored and potentially learned through systematic counterfactual probing of the agent.
Following the generation of tacit knowledge, we model the generation of the counterfactual output $\tilde{Y}_{ij}$ by the agent. This output is conditioned not only on the counterfactual input $\tilde{X}_{ij}$ but also critically on the realized tacit knowledge $h_{ij}$ derived from Equation 8. We define $\tilde{Y}_{ij}$ as the Agent’s generated output when presented with the counterfactual input $\tilde{X}_{ij}$, and $h_{ij}$ acts as the mediating internal state through which $\tilde{X}_{ij}$ is processed to produce $\tilde{Y}_{ij}$.
The generation of $\tilde{Y}_{ij}$ by such agents is often an inherently probabilistic process. Even with a given counterfactual input $\tilde{X}_{ij}$ and a specific internal tacit knowledge state $h_{ij}$, the agent may not produce a single, deterministic output. This inherent stochasticity is captured by the following conditional probability distribution, whose parameters might also be inferred from counterfactual observations:
$\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ (9)
This expression states that $\tilde{Y}_{ij}$ (the counterfactual output) is a random variable whose distribution $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ depends on both the specific tacit knowledge state $h_{ij}$ and the counterfactual input $\tilde{X}_{ij}$. The tacit knowledge $h_{ij}$ here serves as a crucial intermediary. For agents producing sequential outputs, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ might be represented as an autoregressive factorization, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})=\prod_{t}P(\tilde{y}_{ij,t}\mid\tilde{y}_{ij,<t},h_{ij},\tilde{X}_{ij})$.
To determine the overall probability of observing a counterfactual output $\tilde{Y}_{ij}$ from an agent given only the counterfactual input $\tilde{X}_{ij}$—a common scenario when internal states are not directly observable—we must account for the entire hypothesized generative pathway. This involves marginalizing out the unobserved latent tacit variable $\theta_{ij}$ and, by extension, the intermediate tacit knowledge $h_{ij}$. The final generation process is thus expressed as:
[…]
This equation can be derived by considering the dependencies established in Equations 8 and 9.
The derivation proceeds as follows: First, we apply the law of total probability to marginalize over the latent tacit variable $\theta_{ij}$. The probability of $\tilde{Y}_{ij}$ given $\tilde{X}_{ij}$ can be written as an integral over all possible values of $\theta_{ij}$ in its domain $\Theta$:
[…]
Using the definition of conditional probability, $P(A,B\mid C)=P(A\mid B,C)P(B\mid C)$, we can rewrite the integrand $P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})$ as:
$P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})=P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})$
Here, $P(\theta_{ij}\mid\tilde{X}_{ij})$ represents the prior probability distribution of the tacit variable $\theta_{ij}$ given the counterfactual input $\tilde{X}_{ij}$. This term reflects how a given counterfactual input might stochastically lead to different internal configurations or “operational modes” $\theta_{ij}$ within the Agent that are not directly observable. The nature of this prior can also be investigated through counterfactual analysis. Substituting this back, we get:
[…]
The term $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$ is the probability of generating $\tilde{Y}_{ij}$ given both $\theta_{ij}$ and $\tilde{X}_{ij}$. Our model posits a specific structure for this, reflecting a Markov chain-like dependency: $\theta_{ij}\rightarrow h_{ij}\rightarrow\tilde{Y}_{ij}$ (all conditioned on $\tilde{X}_{ij}$ where appropriate). This structure implies two key conditional independence assumptions:
  1. 1.
    Given $\theta_{ij}$, $h_{ij}$ is independent of $\tilde{X}_{ij}$ if all influence of $\tilde{X}_{ij}$ on $h_{ij}$ is mediated through $\theta_{ij}$. More commonly, $P(h_{ij}\mid\theta_{ij})$ is defined as the direct influence of $\theta_{ij}$ on $h_{ij}$.
  2. 2.
    Given $h_{ij}$ and $\tilde{X}_{ij}$, $\tilde{Y}_{ij}$ is independent of $\theta_{ij}$. That is, $P(\tilde{Y}_{ij}\mid h_{ij},\theta_{ij},\tilde{X}_{ij})=P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The tacit knowledge $h_{ij}$ fully mediates the influence of $\theta_{ij}$ on $\tilde{Y}_{ij}$.
Under these assumptions, we can decompose $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$:
[…]
Applying the conditional independence assumptions:
[…]
The form in Equation 10 implies that for a given $\theta_{ij}$ in the outer integral, the relevant $h_{ij}$ is the one generated from that $\theta_{ij}$, effectively collapsing the inner integral. Thus, $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$ is represented by the product $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})$, where $h_{ij}$ is the specific realization tied to $\theta_{ij}$ within the integral. Substituting this construction into Equation 11 directly yields Equation 10:
[…]
This integral sums the probabilities of $\tilde{Y}_{ij}$ occurring via all possible latent tacit variables $\theta_{ij}$. Each path’s contribution is weighted by the prior probability of $\theta_{ij}$, the probability of the corresponding tacit knowledge $h_{ij}$ arising from that $\theta_{ij}$, and finally the probability of the output $\tilde{Y}_{ij}$ given that $h_{ij}$ and the input $\tilde{X}_{ij}$.
Appendix B Implementation Details
Table 6 provides a detailed overview of the datasets, which include three business workflow datasets comprising eleven subsets. Notably, the automotive datasets used for fine-tuning and evaluation are meticulously calibrated by experts from a major automotive manufacturing enterprise.
Table 7 presents a comprehensive comparison of Cochain against other baseline models. Cochain demonstrates superior agent collaboration, characterized by faster inference speeds and lower hallucination rates. More critically, Cochain exhibits decomposability, ensuring that the collaboration process is resilient and does not terminate due to failures in individual components. It is important to note that for CoA, source documents are not included due to the differing task orientation.
Table 6: Dataset Statistics. The table shows detailed information about our three domain-specific business workflow datasets. Seed sources indicate the origin of initial data from which the complete datasets were generated using the Self-Instruct [55] methodology.
Datasets Automotive Pharmaceutical E-commerce
Design Production Supply QC R&D Mfg. Sales Procure. Logistics Operations UX
Total samples 24884 20798 19264 7252 4881 4776 4851 4627 4630 4312 3910
Avg. tokens 958 943 984 870 1656 1601 1623 919 901 917 880
Seed source Industry reports & expert reviewes Clinical databases Platform analytics
The LLMs referenced in Figure 7 are fine-tuned using QLoRA [19], configured with a LoRA rank of 8 and trained for 3 epochs, on two H100 GPUs, with models quantized to 4-bit precision via bitsandbytes. Other open-source models are fine-tuned for 10 epochs using LoRA [25]. The inference times reported in Figure 5 and Table 8 are benchmarked by deploying the open-source models on a single A40 GPU to mitigate network variability.
Table 7: Comparison of different multi-agent collaboration methods
Baseline Collaboration Method Decomposable Reasoning Speed Hallucination Rate Agent
PMC [68] Hierarchical Planning Low Middle Multiple
MedAgents [50] Discussion & Voting Low Middle Multiple
Debate(short) [20] Debate Low Middle Multiple
Debate(long) [20] Debate Low Middle Multiple
CoA [72] Chain Collaboration Low Middle Multiple
Original - High High Single
Cochain Knowledge Fusion High Low Multiple
The appendix provides a more extensive report on experimental results across additional metrics. These include BERTScore-P, BERTScore-R, BLEU-4 [41] for measuring text alignment, METEOR [6] which considers semantically equivalent phrases, ROUGE-2 for assessing content breadth, and ROUGE-L [35] which evaluates deep semantic alignment, particularly for long texts.
Appendix C More Experiments and Analysis
C.1 Efficiency and Cost Analysis
Table 8 presents a comparative analysis of Cochain against other multi-agent systems, focusing on key operational metrics: Inference Time, Token Throughput, and Cost. A noteworthy observation on the Automotive and E-commerce datasets is that Cochain’s average input token count is approximately one-twentieth of that utilized by PMC and CoA. This disparity suggests that these alternative systems expend a considerable volume of tokens on inter-agent communication. In contrast, Cochain innovatively transforms such communication into structured knowledge graphs and prompts tree, a mechanism that effectively curtails operational costs. The overall results compellingly demonstrate that Cochain achieves substantial reductions in inference time, token throughput, and aggregate cost when benchmarked against other multi-agent approaches. Remarkably, the efficiency metrics for Cochain closely approximate those observed for single-agent systems, and its average cost is even lower, underscoring a significant advantage of our proposed framework in terms of computational resource utilization and operational expenditure.
Table 8: Comparative analysis of inference time, token throughput, and cost across three datasets. Inference time is measured using a locally deployed Qwen2-7B model to obviate network latency associated with API calls. Conversely, token throughput and cost metrics are derived using Claude-3.5-Haiku.
Baselines Inference Time (s) Avg.Input Avg.Output Avg.Cost (\$)
Automotive
PMC [68] 241.56 11637.94 6644.22 0.0359
MedAgents [50] 196.53 3901.03 3483.24 0.0170
Debate(short) [20] 187.16 3938.35 4809.27 0.0223
Debate(long) [20] 162.57 4357.87 5278.53 0.0246
CoA [72] 227.21 11428.06 5446.62 0.0309
Original 37.22 93.49 1041.38 0.0042
Cochain 38.77 501.62 541.74 0.0026
Pharmaceutical
PMC [68] 198.49 3796.07 2613.73 0.0135
MedAgents [50] 169.53 3504.85 3504.85 0.0168
Debate(short) [20] 548.66 3372.56 3772.79 0.0178
Debate(long) [20] 574.95 3810.87 4268.74 0.0201
CoA [72] 253.18 5493.94 3450.45 0.0182
Original 60.96 77.28 1091.32 0.0044
Cochain 44.70 533.88 830.34 0.0037
E-commerce
PMC [68] 256.84 11319.74 6452.89 0.0349
MedAgents [50] 181.61 3816.26 3445.34 0.0169
Debate(short) [20] 318.51 4414.23 5660.94 0.0261
Debate(long) [20] 293.42 5438.52 6911.85 0.0321
CoA [72] 195.49 11446.05 5353.24 0.0306
Original 35.83 60.90 1080.92 0.0044
Cochain 37.10 561.80 620.15 0.0029
C.2 More Experiments on Single Agent
Table 9 and Table 10 demonstrate that integrating Cochain markedly enhances single-agent performance across diverse backbone models during two distinct stages of a business workflow. This enhancement is consistently reflected across all evaluated metrics, irrespective of whether the LLMs are in their raw, pre-trained state or have undergone fine-tuning. Notably, a synergistic effect between fine-tuning and the application of Cochain typically yields optimal performance, underscoring Cochain’s robust capability to leverage specialized model knowledge. The broad-spectrum improvements observed across various model metrics highlight the effectiveness and versatility of Cochain in mitigating prior issues of undercollaboration.
Table 9: The performance of our proposed method on single-agent. BERTScore F1 is abbreviated as BS-F, BERTScore Precision as BS-P, and BERTScore Recall as BS-R. We highlight the best and second-best results.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Raw LLM 62.25 64.41 60.54 6.04 8.28 17.52 7.78 15.64
+ Cochain 65.70 68.25 63.50 10.99 13.84 25.50 10.44 16.46
Finetuned LLM 70.68 70.71 70.67 14.31 18.79 35.03 11.97 21.49
+ Cochain 71.46 71.35 71.59 16.50 20.96 38.69 13.79 21.58
Qwen2-7B Raw LLM 71.76 71.79 71.75 17.23 22.30 41.13 12.88 24.00
+ Cochain 75.24 74.90 75.58 22.32 26.51 52.00 19.53 27.94
Finetuned LLM 71.72 71.75 71.70 16.90 21.64 40.77 13.35 24.07
+ Cochain 75.39 75.14 75.65 24.17 28.17 51.22 21.22 28.87
Llama3-8B Raw LLM 68.93 68.38 69.59 7.47 9.14 24.48 10.99 19.28
+ Cochain 73.75 73.36 74.17 19.47 23.52 48.43 17.57 25.10
Finetuned LLM 70.89 70.88 70.91 14.90 19.05 39.83 12.59 22.90
+ Cochain 74.56 74.13 75.01 22.12 26.15 50.59 19.41 27.19
Llama2-13B Raw LLM 60.79 63.16 59.00 5.73 7.96 16.90 7.52 15.43
+ Cochain 67.47 70.20 65.15 13.42 16.88 29.97 13.65 18.57
Finetuned LLM 70.88 70.88 70.90 14.57 19.14 35.87 12.19 21.85
+ Cochain 71.80 72.08 71.58 18.16 22.29 40.33 14.96 23.18
Qwen2.5-14B Raw LLM 71.86 72.03 71.70 9.48 12.16 40.71 11.85 17.86
+ Cochain 75.62 75.56 75.69 12.39 14.71 47.95 17.66 19.55
Finetuned LLM 71.50 71.48 71.53 9.25 11.75 40.16 12.01 17.67
+ Cochain 75.18 75.00 75.38 13.11 15.08 48.35 19.70 20.30
Table 10: The performance of our proposed method. The results of this table are in a different stage, as compared to the stage presented in Table 9. We highlight the best and second-best results.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Raw LLM 61.76 64.72 59.44 6.36 8.69 18.25 8.66 16.17
+ Cochain 63.57 66.31 61.23 9.49 12.37 23.12 8.99 17.46
Finetuned LLM 69.88 69.78 70.00 12.02 17.02 32.59 9.26 20.69
+ Cochain 70.47 70.43 70.55 13.79 18.58 35.44 10.91 20.26
Qwen2-7B Raw LLM 70.34 71.12 70.60 13.74 19.88 35.29 9.47 22.01
+ Cochain 73.48 72.37 74.63 15.77 19.21 50.38 16.25 21.99
Finetuned LLM 70.28 70.08 70.50 13.61 18.95 35.07 9.96 22.38
+ Cochain 74.51 74.30 74.74 21.99 25.76 50.49 20.71 26.00
C.3 More Experiments on Specialized Agent Collaboration
Table 4 presents the results from a distinct experimental configuration designed to further investigate the coordination capabilities of Cochain. Whereas Table 9 assessed a multi-domain model (a single Qwen2-7B model fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), Table 11 exclusively focuses on a setup utilizing multiple agents of the same backbone, each independently fine-tuned for a specific domain. The findings indicate that Cochain effectively facilitates collaboration among these identically architected yet uniquely specialized agents. Performance enhancements are observed both within their respective specialized domains and in the quality of the comprehensive, integrated solutions. This underscores Cochain’s distinct advantage in synergizing specialized knowledge from a team of uniquely trained agents that share a common architectural foundation.
Table 11: Results for one model at particular stages. This means that we use Qwen2-7B for four fine-tunings, and the four fine-tuned models are experimented with at their respective stages. We highlight the fine-tuned results, Cochain application, and improvements()($\uparrow$)( ↑ ).
Backbone Stage Method BS-F BS-P BS-R BLEU-4 GLEU ROUGE-L
Qwen2-7B $S_{1}$ Finetuned 71.72 71.75 71.70 16.90 21.64 24.07
+ Cochain 75.39 75.14 75.65 24.17 28.17 28.87
$\uparrow$ +3.67 +3.39 +3.95 +7.27 +6.53 +4.80
$S_{2}$ Finetuned 70.24 70.47 70.16 13.19 18.29 20.95
+ Cochain 74.23 74.01 74.46 20.05 23.57 24.71
$\uparrow$ +3.99 +3.54 +4.30 +6.86 +5.28 +3.76
$S_{3}$ Finetuned 70.28 70.08 70.50 13.61 18.95 22.38
+ Cochain 74.51 74.30 74.74 21.99 25.76 26.00
$\uparrow$ +4.23 +4.22 +4.24 +8.38 +6.81 +3.62
$S_{4}$ Finetuned 69.40 69.49 69.32 13.51 18.26 21.82
+ Cochain 75.06 74.99 75.15 23.95 27.70 27.49
$\uparrow$ +5.66 +5.50 +5.83 +10.44 +9.44 +5.67
Score 6.4 9.5 +3.1
C.4 More Results from Ablation Studies
On the automotive dataset, we report additional ablation study results focusing on the impact of different backbone models. As indicated in Table 12, “w/o Prompts Tree” markedly impaired performance. Specifically, when applied to the Llama-8B backbone, this ablation led to a 1.97% reduction in the BS-F score, with notable degradations also observed across other evaluation metrics. These findings are consistent with the experimental results presented in Table 5, further underscoring the critical role of the Prompts Tree component in our framework.
Table 12: More results of ablation study. $\mathcal{KG}$ represents the knowledge graph. We highlight the most and second-most efficient modules.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Cochain 71.46 71.35 71.59 16.50 20.96 38.69 13.79 21.58
w/o $\mathcal{KG}$ 71.26 71.14 71.41 16.06 20.53 38.21 13.55 21.50
w/o Causal Chain 71.26 71.13 71.44 15.63 19.59 40.32 14.33 21.16
w/o Prompts Tree 71.06 70.92 71.20 15.20 19.61 36.66 13.07 21.38
Qwen2-7B Cochain 75.39 75.14 75.65 24.17 28.17 51.22 21.22 28.87
w/o $\mathcal{KG}$ 75.21 74.91 75.53 23.56 27.81 49.76 20.15 28.74
w/o Causal Chain 75.20 74.92 75.50 23.68 27.55 51.00 21.14 28.42
w/o Prompts Tree 73.08 73.20 72.98 18.40 22.45 45.43 16.06 24.56
Llama3-8B Cochain 74.56 74.13 75.01 22.12 26.15 50.59 19.41 27.19
w/o $\mathcal{KG}$ 74.27 73.80 74.75 21.45 25.46 49.99 19.09 27.13
w/o Causal Chain 73.63 72.79 74.50 17.34 20.70 49.15 18.28 23.78
w/o Prompts Tree 72.59 72.45 72.75 17.01 21.02 45.00 15.18 23.83
Qwen2.5-14B Cochain 75.18 75.00 75.38 13.11 15.08 48.35 19.70 20.30
w/o $\mathcal{KG}$ 74.83 74.41 75.25 12.49 14.55 47.32 18.39 19.76
w/o Causal Chain 74.71 74.66 74.79 12.28 14.17 47.65 19.30 19.44
w/o Prompts Tree 73.01 73.01 73.02 9.98 12.25 42.72 14.15 18.00
C.5 Integration of Cochain with Baselines
We further investigate the integration of Cochain with existing baseline methods, specifically PMC and MedAgents, through experiments conducted on a pharmaceutical dataset using Claude-3.5-Haiku. As reported in Table 13, the combination of Cochain with PMC results in improvements of 2.62% and 7.55% in the BS-F and BLEU-4 scores, respectively. This outcome further substantiates the efficacy of Cochain in mitigating issues associated with over-collaboration.
Table 13: The improvement of Cochain over other baselines on pharmaceutical datasets.
Baseline BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
PMC [68] 66.51 66.06 66.97 5.17 12.33 20.31 5.48 13.56
PMC [68] + Cochain 69.13 68.49 69.80 12.72 18.48 27.01 6.93 15.82
MedAgents [50] 66.23 66.02 66.48 7.17 14.02 22.59 5.43 14.32
MedAgents [50] + Cochain 67.45 66.91 68.00 11.65 17.80 27.06 6.16 15.74
Appendix D Broader Impacts
Cochain enhances agent efficiency and effectiveness in complex business workflows, potentially increasing productivity and enabling agent systems to tackle more sophisticated multi-step tasks across various industries. By addressing both under-collaboration and over-collaboration, Chain-of-Collaboration contributes to more robust and reliable agent systems. The interpretability offered by our framework, for instance through the causal chain mechanism, can also aid in scrutinizing agent behavior, potentially reducing misuse. Cochain’s design focuses on improving collaboration quality rather than solely on automation. Nevertheless, the broader societal implications of increased automation in workflows, where more tasks are executed by agent systems, warrant ongoing consideration. For the responsible dissemination of our data, which is created using Self-Instruct and consists of generalized, non-sensitive information that does not contain Personally Identifiable Information or private commercial data, we will provide comprehensive documentation, a clear academic research license stipulating ethical use, and a designated point of contact for inquiries.
Appendix E Case Study
E.1 Interpretability Analysis
We investigate the interpretability characteristics of the model after the application of Cochain through a case study. As shown in Table 14, we employ the same color to annotate logically related content visually. The findings reveal that, following the application of Cochain, the model is capable of performing multi-dimensional reasoning based on knowledge graphs, causal inference chains, and contextual cues when addressing user needs, thereby generating outputs that are traceable, reliable, and interpretable. The application of the Cochain method has significantly enhanced the transparency and explainability of the model’s reasoning process, providing empirical evidence for understanding the model’s collaborative decision-making mechanisms.
E.2 Comparative Analysis of Model Outputs
To understand how different baselines handle complex instructions, as shown in Figure 8, we qualitatively analyzed their outputs against a reference answer, focusing on their integration of User Experience(UX) design within the automotive business workflow. Additionally, we present case studies for the pharmaceutical business workflow and the e-commerce business workflow, as shown in Figure 9 and Figure 10.
Cochain: Clear Structure, Focused on UX Integration Process.
Cochain’s response was notably well-structured, using a seven-point list for key UX integration stages, enhancing readability and providing a clear action framework. It comprehensively covered the UX lifecycle—from user research and design principles to interdisciplinary collaboration, technological innovation, continuous iteration, and performance-cost/quality assurance. Its explicit “collaborative perspective of the automotive business workflow” aligned with the reference answer’s holistic view, effectively capturing its spirit by offering a structured, UX-centric methodology.
Comparison of Cochain with Other Model Responses.
Cochain’s primary strength is its sustained focus on the UX design integration process and principles, closely matching the prompt’s core intent and the reference answer. In contrast, PMC and CoA lean towards high-level strategic planning (e.g., PLM platforms, KPIs), emphasizing system management rather than the deep UX integration across workflow stages highlighted by the reference answer. Debate showcases significant depth in UX concepts and cutting-edge technologies. However, its primary focus is on HMI innovation details, rather than systematically integrating UX across broader business workflow processes. Cochain can better balance innovation with this workflow integration. MedAgents share some structural similarities with Cochain, but Cochain is more direct in articulating the end-to-end UX lifecycle, especially “Continuous Iteration and Optimization,” and its business workflow perspective. While other models excel in specific areas (e.g., Debate’s UX tech depth; PMC and CoA’s strategic scope), Cochain demonstrates the closest overall alignment with the reference answer’s advocated UX-integrated process throughout the business workflow and its pragmatic perspective, thus performing excellently in this qualitative assessment.
Table 14: Case study of interpretability. The three distinct modules, namely Knowledge, Causal Chain, and Prompts Tree, are each assigned one of the three different color schemes: red, yellow, and blue. The colors within the modules correspond to those in the output, and logically related content is annotated using the same color scheme.
User need How can cost optimization be achieved in automotive manufacturing?
Knowledge Cost optimization points include technological innovation, quality control, and environmental and energy management. Design optimization, material selection, and supply chain network construction strategies are measures for cost optimization…
Cause chain Stamping is primarily used for component forming and can be integrated with automated and intelligent production lines, intelligent monitoring and optimization systems, as well as smart mold maintenance
Prompts tree Prompt 1: User Need: Cost optimization $\rightarrow$ Design Method: Accelerating prototype iteration $\rightarrow$ Supply Chain Method: Selection of lightweight materials $\rightarrow$ Production Method: Research and analysis of lightweight material performance and cost $\rightarrow$ Quality Inspection Method: ensuring materials meet performance and cost requirements.
Prompt 2: User Need: Cost optimization $\rightarrow$ Design Method: Accelerating prototype iteration $\rightarrow$ Supply Chain Method: Selection of lightweight materials $\rightarrow$ Production Method: Design optimization to enhance material utilization $\rightarrow$ Quality Inspection Method: Testing the strength and stability of the optimized structure.
Output In automotive design, it is essential to consider prototype optimization and innovation in design methodologies. For instance, to address customer-specific vehicle customization, cost control during the design phase can be achieved by accelerating prototype iterations and utilizing lightweight materials. Additionally, the properties and costs of lightweight materials should be taken into account to optimize structural design and enhance material utilization. In supply chain management, cost optimization can also be achieved. This includes strategies for building the supply chain network and optimizing logistics. By consolidating procurement and leveraging economies of scale, procurement costs can be reduced. Collaborative Planning, Forecasting, and Replenishment (CPFR) methods can further optimize inventory costs. In the manufacturing stage, cost optimization can be achieved through the introduction and implementation of automated and intelligent production lines. For example, the use of automated production lines, coupled with intelligent monitoring and optimization systems, can reduce redundant tasks and increase production efficiency. Furthermore, for specific manufacturing processes, such as stamping, material performance costs can be analyzed, and intelligent molds and automatic maintenance systems can be developed to optimize costs. Finally, in the quality inspection stage, ensuring that all materials and products meet the predetermined performance and cost requirements is crucial. Additionally, improving production processes to ensure product processing quality is an important aspect of quality inspection.
Figure 8: Case study on the Automotive dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Figure 9: Case study on the Pharmaceutical dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Figure 10: Case study on the E-commerce dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Appendix F Prompts Template
We provide the prompt templates for all multi-agent baselines presented in the paper, exemplified using the automotive business workflow.
F.1 Cochain
We present a specific overview of the prompts template in the Cochain. As illustrated in Figure 11, the prompt template adopts a modular design comprising three core components: the knowledge module retrieved from the collaborative knowledge graph, the causal chain module, and the business workflow prompts module generated through prompts tree retrieval. In the template, the gray box represents fixed prompt content, while the colored boxes denote dynamic content that adapts based on user needs.
Figure 11: The prompts template of Cochain for final input to LLM.
F.2 PMC
Figure 12: The manager agent prompts template of PMC.
Figure 13: The supervisor agent prompts template of PMC.
Figure 14: The deliverer agent prompts template of PMC.
Figure 15: The executor agent prompts template of PMC.
F.3 MedAgents
Figure 16: The prompt templates for stages 1 through 3 of MedAgents.
Figure 17: The prompt templates for stages 4 and 5 of MedAgents.
F.4 Debate
Figure 18: The prompts template of Debate for final input to LLM.
F.5 CoA
Figure 19: The prompt templates for the worker agent and manager agent of CoA.
Figure 20: The prompt templates for each worker agent on the Automotive dataset.