### v2

# Cochain: Balancing Insufficient and Excessive Collaboration in LLM Agent Workflows

Abstract
Large Language Models (LLMs) have demonstrated impressive performance in executing complex reasoning tasks. Chain-of-thought effectively enhances reasoning capabilities by unlocking the potential of large models, while multi-agent systems provide more comprehensive solutions by integrating the collective intelligence of multiple agents. However, both approaches face significant limitations. Single-agent with chain-of-thought, due to the inherent complexity of designing cross-domain prompts, faces collaboration challenges. Meanwhile, multi-agent systems consume substantial tokens and inevitably dilute the primary problem, which is particularly problematic in business workflow tasks. To address these challenges, we propose Cochain, a collaboration prompting framework that effectively solves the business workflow collaboration problem by combining knowledge and prompts at a reduced cost. Specifically, we construct an integrated knowledge graph that incorporates knowledge from multiple stages. Furthermore, by maintaining and retrieving a prompts tree, we can obtain prompt information relevant to other stages of the business workflow. We perform extensive evaluations of Cochain across multiple datasets, demonstrating that Cochain outperforms all baselines in both prompt engineering and multi-agent LLMs. Additionally, expert evaluation results indicate that the use of a small model in combination with Cochain outperforms GPT-4.
**footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Introduction
Figure 1: Collaboration has a Goldilocks region. With Insufficient Collaboration, a system overlooks cross-stage constraints and dependencies. With Excessive Collaboration, interaction injects peripheral details that dilute the core decision and increase token cost. Cochain targets effective collaboration by enforcing stage constraints while keeping the solution focused. We illustrate this with a car manufacturing query, where a single agent misses manufacturing and supplier constraints, and excessive discussion dilutes focus on issues such as inventory rollout and price volatility.
Large Language Models (LLMs) have achieved strong performance on language understanding and complex reasoning tasks Touvron et al. (2023a); Bai et al. (2023); GLM et al. (2024); Bi et al. (2024). Much of this progress is unlocked at inference time. Prompting techniques such as Chain-of-Thought encourage step-by-step reasoning and often improve depth of reasoning Wei et al. (2022); Besta et al. (2024). In parallel, multi-agent frameworks extend a single model into a team of specialized roles and interaction protocols, aiming to provide broader coverage across domains such as healthcare, education, law, and finance Tang et al. (2024); Dan et al. (2023); Cui et al. (2024); Yang et al. (2023). However, a practical question remains: How much collaboration is actually needed for high-quality decision making?
This question is especially salient in business workflow tasks. We define a Business Workflow as a complex process consisting of many interconnected stages. The stages are specialized, and they strongly depend on each other. A decision made in one stage requires more than local expertise. It must anticipate and incorporate knowledge and constraints from other stages. For example, in an automotive manufacturing workflow, design and production choices must align with what the supply chain can actually provide. In this setting, we observe two opposite failure modes (Figure 1). Insufficient Collaboration arises when the overall system does not effectively integrate downstream constraints into the decision process, so solutions may look locally reasonable but violate workflow level requirements Sahoo et al. (2024). Excessive Collaboration arises when many agents participate directly and simultaneously in each decision, making interaction expensive and unfocused, so non critical information overwhelms the main issue and the system may prefer consensus-like answers over accurate ones Du et al. (2024); Zhang et al. (2024b). Existing research on multi-agent systems primarily focuses on maximizing collaboration, extensively exploring how multi-agent systems improve decision quality Zhang et al. (2025b), make safe decisions Piatti et al. (2024), and solve complex problems Li et al. (2025). In contrast, the negative effects of excessive collaboration have received insufficient attention, particularly in multi-stage business workflows Lei et al. (2025). To our best knowledge, no prior work has systematically analyzed and defined this Excessive Collaboration phenomenon in workflow settings, nor proposed targeted solutions to mitigate it. We therefore seek effective collaboration that is stage aware and focused, without paying the full cost of direct multi-agent chatting.
To this end, we propose Cochain, a chain-of-collaboration prompting framework for business workflows. Cochain replaces token-intensive direct interaction with retrieval of reusable collaboration artifacts distilled from agents and data (Figure 2). It constructs a collaborative knowledge graph that fuses explicit knowledge from the original datasets with tacit knowledge elicited from specialized agents via counterfactual questioning. To mitigate the influence of irrelevant knowledge Zhou et al. (2024), retrieved evidence is organized into compact causal chains that connect stage-specific information through shared bridge entities, supporting cross-stage reasoning while filtering noise. Cochain also builds a prompts tree that distills reusable, solution-oriented prompts from agent responses and organizes them across workflow stages. Inspired by how the human brain integrates fragmented knowledge into coherent thought Courellis et al. (2024), the prompt tree is designed to provide structured cross-stage guidance: querying the tree yields a prompt chain tailored to the input, guiding the backbone model through stage-specific reasoning in a consistent order. At inference time, Cochain retrieves stage-relevant knowledge, a causal chain from the knowledge graph, and a prompt chain from the prompt tree, then composes them into a structured prompt for the backbone LLM. This produces workflow-aware yet focused answers, balancing insufficient collaboration and excessive collaboration.
The main contributions are summarized as follows:
  • Effective collaboration in business workflows. We identify two opposite collaboration failures in business workflow tasks, insufficient collaboration that misses cross-stage constraints and excessive collaboration that dilutes the core decision with excessive interaction.
  • Cochain framework. We propose Cochain, which enables stage-aware collaboration via reusable artifacts rather than token-heavy interactions. Cochain combines a collaborative knowledge graph with causal-chain retrieval and a prompts tree that composes cross-stage prompt chains for coherent multi-stage reasoning.
  • Empirical evidence and expert support. Experiments across benchmarks and backbones show that Cochain improves quality with efficient inference. Expert evaluation further suggests a smaller model with Cochain can outperform a larger one. We will release our collaborative knowledge graph to support future research.
Figure 2: Overview of Cochain. Offline, we construct a Collaborative Knowledge Graph (CKG) by combining explicit triples from workflow data with tacit triples elicited via counterfactual querying and iterative verification, and build a Prompts Tree through cross-stage prompt distillation. Online, Cochain retrieves stage-relevant knowledge from the CKG, composes causal chains, and leverages the Prompts Tree to propagate necessary constraints across stages, mitigating insufficient collaboration and excessive collaboration.
2 Related Work
2.1 Knowledge Graph Augmented LLMs
LLM reasoning capabilities are critical for high-quality responses Jain et al. (2024). Knowledge graphs, with their structured, explicit, and interpretable nature Zhou et al. (2025), provide clear representations and transparent reasoning paths, guiding LLMs toward deeper reasoning and mitigating limitations Wang et al. (2024). For domain-specific LLMs, knowledge graphs enable continuous knowledge updates to accommodate evolving information Lavrinovics et al. (2025). Applications include MindMap Wen et al. (2024) in medicine and ChatLAW Cui et al. (2024) in law. Business workflows pose an additional challenge of multi-stage, cross-domain knowledge fusion, where irrelevant knowledge can easily overwhelm the core decision and degrade performance Ouyang et al. (2022). To address this, Cochain organizes retrieved evidence into causal chains, improving cross-stage understanding while filtering noise.
2.2 Prompt Engineering
Prompt engineering guides LLMs to maximize their potential via designed prompts Sahoo et al. (2024). It enables model adaptation to new tasks through in-context learning and instruction following Brown et al. (2020); Li et al. (2023). Well-designed prompts are shown to enhance performance, particularly on complex tasks Wei et al. (2022); Wang et al. (2023); Yao et al. (2023). CoT Wei et al. (2022) enables models to generate intermediate reasoning steps via simple prompts. ToT Yao et al. (2024) explores coherent text units as intermediate problem-solving steps. CPO Zhang et al. (2024a) leverages non-optimal reasoning paths from tree search to speed up inference. However, these methods are typically designed for a single decision context and do not explicitly maintain constraint awareness across stages or support cross-domain coordination in business workflows. Cochain complements this line by distilling cross-stage prompts from agent responses and organizing them into prompts tree for retrieval.
2.3 Collaboration of LLM Agents
Many studies in LLM research solve problems through multi-agent collaboration Du et al. (2024); Zhao et al. (2024). Common paradigms include discussion Tang et al. (2024) and debate Du et al. (2024) to enhance reasoning, hierarchical structures with specialized roles Zhang et al. (2025a), and sequential or tree-structured architectures Zhang et al. (2024b); Zhao et al. (2024). These have proven effective in domains like medicine Tang et al. (2024), long-text processing Zhang et al. (2024b); Zhao et al. (2024), social simulation Wei et al. (2024), and code intelligence Huang et al. (2023). Most existing systems rely on richer interaction and repeated message exchanges, which increases token cost and can dilute the core decision in multi-stage workflows when peripheral opinions accumulate. Cochain instead uses reusable collaboration artifacts to deliver focused, stage-aware collaboration while keeping inference lightweight. To our best knowledge, no prior work has systematically defined Excessive Collaboration in this setting or proposed a targeted mitigation mechanism.
Backbone Baseline Automotive Pharmaceutical E-commerce
BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L
Pangu-1B PMC ${\text{72.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{12.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{22.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{75.67}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{20.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{29.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{73.96}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{11.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.40}}$ ${\text{19.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$
MedAgents ${\text{69.95}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.22}}$ ${\text{10.40}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{20.22}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{72.82}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{17.72}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{27.12}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.37}}$ ${\text{74.63}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.16}}$ ${\text{21.22}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.34}}$ ${\text{28.99}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$
Debate(short) ${\text{71.36}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.11}}$ ${\text{10.72}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{21.45}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{75.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$ ${\text{21.87}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{30.11}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{75.73}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.33}}$ ${\text{28.05}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$ ${\text{31.05}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$
Debate(long) ${\text{70.59}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.20}}$ ${\text{10.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{20.85}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{73.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.37}}$ ${\text{18.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{27.61}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.15}}$ ${\text{75.38}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{16.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.34}}$ ${\text{25.27}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$
CoA ${\text{69.61}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{8.54}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.47}}$ ${\text{18.60}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{73.82}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.33}}$ ${\text{18.79}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{28.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.33}}$ ${\text{74.75}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{17.32}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$ ${\text{24.10}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$
Cochain ${\text{74.35}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{15.99}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{26.01}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.03}}$ ${\text{76.40}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{22.10}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{30.71}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{77.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.04}}$ ${\text{23.37}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{31.64}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.11}}$
Qwen2-7B PMC ${\text{65.86}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.03}}$ ${\text{13.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{18.35}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{66.97}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.79}}$ ${\text{16.25}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.58}}$ ${\text{24.75}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.62}}$ ${\text{71.02}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{24.18}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.37}}$ ${\text{29.93}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.45}}$
MedAgents ${\text{65.16}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{12.37}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{18.33}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{65.84}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.15}}$ ${\text{10.68}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.36}}$ ${\text{20.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.50}}$ ${\text{69.45}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{17.42}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{33.38}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$
Debate(short) ${\text{65.30}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{12.42}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.37}}$ ${\text{17.63}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{70.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$ ${\text{16.04}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{24.12}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{73.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$ ${\text{20.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.32}}$ ${\text{28.27}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$
Debate(long) ${\text{65.61}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{12.85}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.47}}$ ${\text{18.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{60.57}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$ ${\text{10.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{11.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.52}}$ ${\text{72.41}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.60}}$ ${\text{20.82}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.59}}$ ${\text{28.46}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.53}}$
CoA ${\text{70.70}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.53}}$ ${\text{21.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$ ${\text{22.65}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.68}}$ ${\text{76.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{30.45}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.42}}$ ${\text{34.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{79.24}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{38.60}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{39.14}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.62}}$
Cochain ${\text{75.05}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{27.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{28.43}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{78.50}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.04}}$ ${\text{34.48}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$ ${\text{35.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{80.22}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.16}}$ ${\text{37.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.38}}$ ${\text{40.84}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.41}}$
DeepSeek-R1-7B PMC ${\text{65.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{10.75}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{15.78}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{68.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{22.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{28.07}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.20}}$ ${\text{66.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$ ${\text{15.80}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{23.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$
MedAgents ${\text{65.58}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.60}}$ ${\text{8.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.55}}$ ${\text{14.08}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.59}}$ ${\text{70.35}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{24.47}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$ ${\text{30.90}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{66.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{13.67}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{20.31}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.60}}$
Debate(short) ${\text{65.93}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{11.41}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{17.33}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.52}}$ ${\text{71.46}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{27.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.27}}$ ${\text{30.34}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$ ${\text{70.71}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{23.01}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.22}}$ ${\text{27.02}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$
Debate(long) ${\text{66.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.67}}$ ${\text{12.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{17.70}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{71.54}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.42}}$ ${\text{27.34}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{30.04}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{70.48}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{23.07}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{27.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.44}}$
CoA ${\text{70.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{13.27}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.58}}$ ${\text{17.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.58}}$ ${\text{77.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.15}}$ ${\text{38.18}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{35.53}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$ ${\text{80.43}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.32}}$ ${\text{43.80}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{42.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$
Cochain ${\text{73.20}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.02}}$ ${\text{21.45}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{23.89}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{79.11}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$ ${\text{38.27}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{37.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$ ${\text{82.68}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{50.19}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{47.07}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$
Claude-3.5-haiku PMC ${\text{64.06}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.04}}$ ${\text{9.73}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.56}}$ ${\text{10.12}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{66.18}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.16}}$ ${\text{13.04}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.36}}$ ${\text{13.91}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.18}}$ ${\text{66.85}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{15.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.60}}$ ${\text{15.88}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.21}}$
MedAgents ${\text{64.38}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{10.77}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.56}}$ ${\text{11.02}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{66.01}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{14.60}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.30}}$ ${\text{14.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{68.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{17.58}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{16.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$
Debate(short) ${\text{64.36}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{11.21}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{13.72}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.33}}$ ${\text{69.11}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.11}}$ ${\text{14.53}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{16.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$ ${\text{68.41}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.43}}$ ${\text{19.97}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.32}}$ ${\text{18.37}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$
Debate(long) ${\text{65.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{11.79}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.49}}$ ${\text{12.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{67.68}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.10}}$ ${\text{15.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.31}}$ ${\text{15.13}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.32}}$ ${\text{67.49}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.67}}$ ${\text{17.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.42}}$ ${\text{18.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.56}}$
CoA ${\text{63.78}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.26}}$ ${\text{9.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.29}}$ ${\text{10.28}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.20}}$ ${\text{66.00}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.39}}$ ${\text{13.33}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.22}}$ ${\text{14.25}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.36}}$ ${\text{67.03}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$ ${\text{15.02}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.33}}$ ${\text{15.12}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.35}}$
Cochain ${\text{69.66}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.03}}$ ${\text{18.62}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.06}}$ ${\text{16.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.24}}$ ${\text{72.83}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.06}}$ ${\text{23.47}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{19.25}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.22}}$ ${\text{71.56}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.12}}$ ${\text{20.11}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.13}}$ ${\text{20.65}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$
Table 1: Performance comparison of different model backbones and multi-agent baselines across three business domains. We report the mean and standard error (SE) over five experiments. BERTScore F1 is abbreviated as BS-F. We highlight the best and second-best results.
3 Methods
3.1 Collaborative Knowledge Graph
Building a collaborative knowledge graph requires abundant stage-specific knowledge. We obtain it from the training sets of the workflow datasets, denoted as $D_{i}=\{(X_{ij},Y_{ij})\}_{j=1}^{n_{i}}$, where $i\in\{1,\dots,N\}$ indexes workflow stages, $N$ is the number of stages, and $j\in\{1,\dots,n_{i}\}$ indexes the $n_{i}$ training instances in stage $i$. Here $X_{ij}\in X_{i}$ is the input and $Y_{ij}\in Y_{i}$ is the corresponding output. After data cleaning and triplet extraction, we construct an explicit knowledge graph $\mathcal{KG}_{\text{explicit}}$:
$\mathcal{KG}_{\text{explicit}}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{n_{i}}\text{ExtractTriples}(X_{ij},Y_{ij})$ (1)
Business workflows also contain tacit knowledge that is rarely stated in the original datasets, such as feasibility constraints, implementation caveats, and stage-specific heuristics. Since such considerations are not explicitly available, we elicit tacit workflow knowledge from the agent’s responses by probing the agent under counterfactual perturbations and attributing response variations to unobserved internal considerations.
For each sample $(X_{ij},Y_{ij})$, we generate a counterfactual input $\tilde{X}_{ij}$ by editing $X_{ij}$ to perturb one salient condition while preserving intent, using templates for causal, adversarial, substitution, extreme, and backward-causal variants:
$\tilde{X}_{ij}=\text{GenerateCounterfactual}(X_{ij})$ (2)
In our setting, a vertical domain agent is the stage-specialized LLM assigned to stage $i$ (fine-tuned on stage-$i$ data for open-source backbones). We query this agent with $\tilde{X}_{ij}$ to obtain the counterfactual output $\tilde{Y}_{ij}$. To capture the unobserved considerations that shape how the agent resolves $\tilde{X}_{ij}$, we introduce a latent variable $\theta_{ij}$ and an intermediate latent state $h_{ij}$ influenced by $\theta_{ij}$, where $h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$, and $\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The resulting generation process can be written as:
$\begin{split}P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})&=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})\\ &\quad\cdot P(h_{ij}\mid\theta_{ij})\,P(\theta_{ij}\mid\tilde{X}_{ij})\,d\theta_{ij}\end{split}$ (3)
Here, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ is the likelihood of generating the counterfactual output given $h_{ij}$ and $\tilde{X}_{ij}$, $P(h_{ij}\mid\theta_{ij})$ describes how tacit considerations shape the latent state, and $P(\theta_{ij}\mid\tilde{X}_{ij})$ is a prior over tacit considerations conditioned on the counterfactual input.
To improve reliability, each counterfactual answer is evaluated by a general-purpose LLM and assigned one of three labels: reasonable, ambiguous, or unreasonable. The evaluator also provides concise feedback and revision suggestions, which the vertical agent uses to regenerate a response. This loop repeats until a reasonable answer is obtained. During knowledge distillation, we apply causal enhancement by emphasizing causal cues such as “depends on”, “relies on”, and “applies to”, which helps preserve cross-stage dependency signals. Using the counterfactual input-output pairs $(\tilde{X}_{ij},\tilde{Y}_{ij})$ produced by the refinement loop, we extract tacit triples and build a tacit knowledge graph $\mathcal{KG}_{\text{tacit}}$:
$\mathcal{KG}_{\text{tacit}}=\bigcup_{i=1}^{N}\bigcup_{j=1}^{n_{i}}\text{ExtractTriples}(\tilde{X}_{ij},\tilde{Y}_{ij},\theta_{ij})$ (4)
Finally, we integrate explicit and tacit knowledge to form the collaborative knowledge graph $\mathcal{KG}$, which captures both stage-specific knowledge and inter-stage connectivity:
$\mathcal{CKG}=\mathcal{KG}_{\text{explicit}}\cup\mathcal{KG}_{\text{tacit}}$ (5)
3.2 Causal Chain
To help the agent organize cross-stage dependencies instead of merely listing retrieved facts, we introduce a causal chain mechanism. Given a user request, we decompose it and extract keywords, which are then used to retrieve relevant nodes from the collaborative knowledge graph.
Following Figure 2, we use a pre-trained text encoder Reimers and Gurevych (2019) to embed each keyword and candidate knowledge node. For the embedding $v_{k}$ of keyword $k$ and the embedding $v_{n}$ of node $n$, we compute cosine similarity:
$\text{Sim}(v_{k},v_{n})=\frac{v_{k}\cdot v_{n}}{\|v_{k}\|\|v_{n}\|}$ (6)
With a similarity threshold $\delta$, we select nodes that are semantically related to the query. For each selected node, we expand one-hop neighbors to introduce closely associated evidence and form a local triple-based structure. To connect knowledge across workflow stages, we further perform multi-hop exploration through bridge entities. Let the workflow stages be $\mathcal{S}=\{s_{1},s_{2},s_{3},\dots\}$. For a high-similarity node $A$ belonging to stage $s_{1}$, we expand a neighbor $B$ that appears in both $s_{1}$ and some other stage $s_{i}$ with $i\neq 1$. We then use $B$ as a bridge to reach a node $C$ in stage $s_{i}$, forming a cross-stage causal chain:
$\text{CausalChain}(A,C)=A\xrightarrow{B}C$ (7)
The causal chain provides a compact, stage-connected context that highlights how constraints and decisions propagate across stages, improving coherence and feasibility in the final response.
Backbone Method BS-F GLEU ROUGE-L
GPT-3.5-turbo + IO ${\text{72.84}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{15.09}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{21.98}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.23}}$
+ CoT ${\text{71.61}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{11.15}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{17.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.16}}$
+ ToT ${\text{66.69}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.05}}$ ${\text{10.08}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.04}}$ ${\text{17.74}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$
+ Cochain ${\text{76.94}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.03}}$ ${\text{28.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.12}}$ ${\text{30.65}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.09}}$
GPT-4o + IO ${\text{71.92}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.08}}$ ${\text{23.05}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{23.79}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$
+ CoT ${\text{72.57}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.15}}$ ${\text{23.76}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{24.15}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.28}}$
+ ToT ${\text{66.23}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.17}}$ ${\text{12.51}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$ ${\text{17.52}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.19}}$
+ Cochain ${\text{74.46}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.07}}$ ${\text{25.16}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.25}}$ ${\text{25.07}}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\,0.14}}$
Table 2: The performance of GPT-3.5-turbo and GPT-4o when using CoT, ToT, and Cochain.
Baseline Backbone
DeepSeek-V3.1 Claude-3.5-haiku
PMC 0.12 0.29
MedAgents 0.07 0.26
Debate 0.32 0.31
CoA 0.27 0.37
Cochain 0.33 0.38
Table 3: Accuracy comparison on Auto-SLURP benchmark.
3.3 Prompts Tree
In business workflow tasks, effective collaboration depends on whether stage-specific reasoning can be carried forward in a consistent and reusable manner. We propose a prompt distillation procedure that offline constructs a Prompts Tree, which stores reusable, solution-oriented prompts and explicitly encodes cross-stage dependencies. We model the Prompts Tree as a rooted tree $T=(V,E)$. Each node $v\in V$ is associated with a stage label $s(v)\in\mathcal{S}$, a distilled prompt $p(v)$, and its supporting evidence (the Q&A from which the prompt is distilled). Each directed edge $(u\!\rightarrow\!v)\in E$ represents a dependency where prompt $p(u)$ from an upstream stage induces a downstream question template, whose answer leads to prompt $p(v)$ at a later stage. Querying $T$ returns a prompt chain along a root-to-leaf path, providing structured, cross-stage guidance during inference.
Tree construction. We start from a seed Q&A at Stage 1. Given the Stage 1 answer, the Stage 1 agent distills a set of candidate prompts that are solution-oriented and constraint-aware, focusing on actionable decisions, feasibility conditions, and implementation requirements. Each candidate prompt is then self-evaluated for utility and executability, and we retain the top $m$ prompts. For each retained prompt, we generate a downstream question template that asks how to implement, validate, or satisfy this prompt in the next stage. The next-stage agent answers these questions, producing new Q&A pairs that serve as seeds for prompt distillation at the next stage. This distill–self-evaluate–generate-next-question–expand loop repeats across stages, causing the tree to grow as upstream prompts branch into multiple downstream realizations.
Node and edge semantics. The user need is treated as the root node. Prompts distilled at Stage 1 become children of the root. For a node $u$ at stage $s_{k}$, each outgoing edge corresponds to a generated question template $q_{u\rightarrow}$ targeting stage $s_{k+1}$. The resulting answer yields a child node $v$ at stage $s_{k+1}$ with distilled prompt $p(v)$. This construction explicitly records cross-stage dependencies, since each child prompt is conditioned on satisfying or operationalizing its parent prompt under downstream constraints.
Retrieval and usage. Inspired by how the human brain integrates fragmented knowledge into coherent thought Courellis et al. (2024), the Prompts Tree organizes fragmented stage-specific guidance into an ordered chain. At inference time, we retrieve a prompt chain $\pi=(v_{0}\!\rightarrow v_{1}\!\rightarrow\cdots\!\rightarrow v_{L})$ that best matches the input query, and use the prompts $\{p(v_{\ell})\}_{\ell=0}^{L}$ as structured guidance so the backbone LLM reasons through the workflow stages in a consistent order. This avoids brittle manual prompt construction while improving cross-stage coordination with low token overhead.
4 Experiments
We study when collaboration helps and when it hurts in multi-stage business workflows, and evaluate whether Cochain provides a practical middle ground via reusable collaboration artifacts. Accordingly, we benchmark overall performance against strong baselines, analyze key design factors and hyperparameters, compare jointly fine-tuned and stage-wise specialized deployment regimes, and report efficiency and cost.
4.1 Experimental Setup
Datasets.
We evaluate Cochain on two benchmarks across four datasets: Auto-SLURP Shen and Shen (2025), which assesses end-to-end multi-stage workflows, and MSCoRe Lei et al. (2025), which focuses on multi-stage reasoning and collaboration. The automotive workflow test set is reviewed and refined by domain experts from a major automobile manufacturer to ensure practical industry relevance.
Metrics.
We report BERTScore Zhang et al. (2019), GLEU Wu et al. (2016), and ROUGE-L Lin (2004) to measure semantic similarity and generation quality between model outputs and reference answers.
LLMs.
We use 17 backbone LLMs, covering both proprietary and open-source models. We compare reasoning-capable models (e.g., DeepSeek-R1-7B DeepSeek-AI (2025)) with standard models (e.g., Pangu-1B Chen et al. (2025), Qwen2-7B 29). We also study using identical backbones across stages versus mixed backbones (Llama2-7B Touvron et al. (2023b), GLM4-9B, Qwen2-7B, DeepSeek-7B) for stage specialization. For open-source backbones, we perform stage-wise fine-tuning on MSCoRe: the agent at stage $k$ is fine-tuned using only the training split of stage-$k$ data, without using any validation/test instances. Proprietary models include GPT-3.5-turbo, GPT-4o Achiam et al. (2023), DeepSeek-V3.1 DeepSeek-AI (2024), and Claude-3.5-haiku.
Figure 3: Expert evaluation and GPT-4 scoring results. Panel (a) compares Qwen2-7B augmented with Cochain against GPT-4 and expert-written references (100 cases). In panel (b), the asterisk indicates augmentation with Cochain.
Baselines.
Our baseline selection encompasses the predominant multi-agent collaboration methodologies. Specifically, PMC Zhang et al. (2025a) implements multi-agent collaboration through hierarchical planning. MedAgents Tang et al. (2024) facilitates collaborative decision-making via discussion and voting mechanisms. Debate Du et al. (2024) enables collaboration through argumentative discourse. CoA Zhang et al. (2024b) establishes chain-based collaboration. Besides, we compare the performance of single-agent prompt engineering methods, namely CoT Wei et al. (2022) and ToT Yao et al. (2024).
Skip Stage BS-F GLEU ROUGE-L
zero / 75.39 28.17 28.87
one $S_{1}\rightarrow S_{3}\rightarrow S_{5}$ 74.00 24.66 25.18
$S_{2}\rightarrow S_{4}$ 74.52 25.57 26.18
$S_{3}\rightarrow S_{5}$ 74.18 25.45 25.95
two $S_{1}\rightarrow S_{3}$ 74.27 25.26 25.74
$S_{2}\rightarrow S_{4}$ 74.07 25.00 25.82
Table 4: Result of different skip stages on evaluation metrics.
4.2 Result
Overall Results.
Table 1 summarizes performance across three business domains and four backbones. The results support our central claim that effective workflow collaboration requires balancing Insufficient and Excessive Collaboration. Compared with multi-agent baselines that rely on rich interaction (often leading to excessive collaboration and diluted decisions), Cochain consistently achieves higher quality across domains. Meanwhile, Cochain also improves substantially on a small backbone, indicating that it mitigates insufficient collaboration by injecting stage-aware constraints and dependencies through reusable artifacts rather than relying on stronger backbones alone. For instance, with Pangu-1B on Automotive, Cochain raises BS-F from 72.28 (PMC) to 74.35 and improves ROUGE-L from 22.89 to 26.01, suggesting that the framework helps the agent stay constraint-aware while remaining focused on the core decision.
Benchmark Transfer on Auto-SLURP.
Table 3 confirms the advantage on an end-to-end workflow benchmark. Cochain achieves the highest accuracy under both DeepSeek-V3.1 and Claude-3.5-haiku backbones, reaching 0.33 and 0.38, respectively. This result supports that Cochain generalizes beyond MSCoRe-style text similarity metrics and remains effective when collaboration must be executed across workflow stages.
Comparison to Prompting Baselines.
Table 2 compares Cochain with IO, CoT, and ToT on proprietary backbones. While CoT and ToT enhance reasoning in a single-agent setting, they provide limited support for carrying stage constraints across a workflow. Cochain yields stronger generation quality under both GPT-3.5-turbo and GPT-4o. On GPT-3.5-turbo, Cochain increases BS-F from 72.84 (IO) to 76.94 and improves ROUGE-L from 21.98 to 30.65, demonstrating that the benefits go beyond deeper single-path reasoning.
Expert Evaluation and GPT-4 Scoring.
Figure 3 provides external validation beyond automatic metrics. In panel (a), domain experts judge Cochain-augmented answers to be better in 40 cases and worse in 9 cases, with 51 cases considered similar. Panel (b) shows higher GPT-4 scores for settings augmented with Cochain, and the fine-tuned augmented setting reaches near-ceiling scores, corroborating the expert preference trend.
Figure 4: Hyperparameter influence of prompt budget and retrieval size. (a) Fix top-$n$ retrieved nodes and vary prompt count. (b) Fix prompt count and vary top-$n$.
Figure 5: Comparative analysis of metrics scaling with increasing model size from DeepSeek-R1 7B to 70B.
Figure 6: Stage-wise gains of Cochain over fine-tuning on the Automotive workflow ($\Delta=\text{+Cochain}-\text{Finetuned}$). (a) Jointly fine-tuned backbone (Qwen2-7B). (b) Stage-wise fine-tuned backbones. Bars show improvements in BS-F, GLEU, and ROUGE-L. The box reports the overall GPT-4 score gain of the integrated solution.
Backbone Method Automotive Pharmaceutical E-commerce
BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L BS-F GLEU ROUGE-L
Qwen2-7B Cochain 75.39 28.17 28.87 78.48 33.98 35.03 79.91 37.08 40.06
w/o $\mathcal{CKG}$ 75.21 27.81 28.74 76.34 25.92 28.09 76.72 32.40 32.74
w/o Causal Chain 75.20 27.55 28.42 76.20 25.38 27.76 76.86 33.30 32.93
w/o Prompts Tree 73.08 22.45 24.56 76.19 26.84 28.91 76.59 31.83 32.38
DeepSeek-R1-7B Cochain 73.19 21.73 24.01 78.97 37.94 36.75 82.68 50.30 47.24
w/o $\mathcal{CKG}$ 72.41 19.71 22.63 78.22 37.33 36.40 77.36 36.54 35.85
w/o Causal Chain 72.67 19.96 22.86 78.41 37.36 36.74 77.60 36.30 35.74
w/o Prompts Tree 72.27 16.54 20.60 76.96 37.98 35.61 79.02 41.03 38.84
Claude-3.5-haiku Cochain 69.61 18.73 17.30 72.87 23.59 19.67 71.36 20.15 20.78
w/o $\mathcal{CKG}$ 68.98 16.63 16.31 70.93 16.42 18.45 71.13 18.85 20.22
w/o Causal Chain 69.33 17.31 16.65 71.33 17.43 18.78 71.23 18.89 20.45
w/o Prompts Tree 65.79 12.84 13.71 70.85 16.24 18.31 69.76 15.55 18.59
Table 5: Result of ablation study. We highlight the largest and second-largest performance drops caused by removing each component.
4.3 Key Design Factors
Ablation Study.
To validate Cochain’s effectiveness, we conduct ablation experiments with three variants: (1) w/o Knowledge Graph; (2) w/o Causal Chain; and (3) w/o Prompts Tree. As shown in Table 5, all variants exhibit performance degradation across all metrics, confirming each component’s importance. Notably, removing the prompts tree causes the most significant drop, highlighting its critical role in collaboration. The knowledge graph and causal chain components also prove essential, demonstrating their complementary functions in the reasoning process.
Hyperparameter Influence.
We vary two hyperparameters: the prompt count and the retrieval nodes top-$n$ from the knowledge graph. Figure 4 shows an inverted U-shape when fixing top-$n$ and increasing prompt count, indicating an optimal prompt budget rather than monotonic gains. When fixing the prompt budget and varying top-$n$, the trend is not monotonic. Overall, performance is more sensitive to prompt budgeting than simply retrieving more nodes, consistent with Table 5 where Prompts Tree yields the largest contribution.
Stage Dependency.
Table 4 shows that skipping any stage consistently degrades all metrics, indicating that each stage contributes non-redundant constraints for end-to-end workflow reasoning.
Figure 7: Cost comparison (Tokens, API Fees, Time) across baselines on pharmaceutical dataset. Cochain’s total costs include the one-time Prompts Tree setup and costs for varying evaluation queries.
4.4 Cross-Domain and Specialized Settings
We evaluate Cochain on the Automotive workflow under two regimes: a jointly fine-tuned backbone trained on the union of all stage data, and stage-wise fine-tuned backbones specialized per stage. Figure 6 reports stage-wise gains over the corresponding fine-tuned baselines, where $\Delta=\text{+Cochain}-\text{Finetuned}$. With joint fine-tuning (Figure 6(a)), Cochain improves all stages and metrics, with larger gains in later stages where downstream constraints are more tightly coupled with upstream decisions, indicating better constraint propagation without direct, token-intensive interaction. With stage-wise fine-tuning (Figure 6(b)), Cochain consistently boosts each specialized backbone, showing it can coordinate domain-expert agents into a coherent workflow solution. The integrated four-stage outputs also receive higher GPT-4 scores in both regimes, supporting end-to-end quality improvements beyond stage-local changes.
4.5 Efficiency and Cost Analysis
Figure 7 breaks down efficiency into one-time setup overhead and per-query running cost. In contrast to interaction-heavy baselines, Cochain shifts most computation to an offline setup stage and keeps inference lightweight. This yields consistent savings across budgets: for 1000 evaluation queries, Cochain consumes only 1.98M tokens in total, compared with 8.94M for CoA and 8.08M for Debate(long). The same trend holds in monetary cost, where Cochain stays at \$5.80 versus \$18.20 and \$20.12, and in wall-clock time, where Cochain finishes in 79,923s versus 253,184s and 574,949s. While the setup stage introduces an upfront cost, it is paid once and then reused, so the marginal cost grows slowly with more queries and the total advantage widens at scale. This shows that Cochain delivers high-quality solutions at low cost, effectively alleviating excessive collaboration.
5 Conclusion
In this paper, we propose Cochain, a collaboration prompting framework that balances Insufficient Collaboration and Excessive Collaboration in LLM agent workflows. Cochain bridges single-agent depth and multi-agent breadth via reusable collaboration artifacts, achieving low operational cost, fast inference, strong interpretability, and robust end-to-end reasoning. It builds a Prompts Tree to distill and propagate stage-wise constraints, enabling coherent multi-stage solutions without token-heavy interactions. Cochain also constructs a cross-stage Collaborative Knowledge Graph with tacit knowledge elicited through counterfactual querying, which strengthens relational reasoning and reduces hallucinations. In addition, the causal chain mechanism exposes explicit dependency traces, improving transparency and controllability. Extensive experiments on specialized and cross-domain settings show consistent improvements in scenarios affected by both collaboration extremes, and expert evaluation further indicates that smaller models equipped with Cochain can outperform GPT-4 on workflow tasks.
Appendix A Algorithms and Examples to Better Understand Cochain
Algorithm 1 Collaborative Knowledge Graph Construction
1:Datasets $D=\{D_{1},\dots,D_{N}\}$; LLM agents $A=\{A_{1},\dots,A_{N}\}$; Evaluation LLM $E$; Stage Labels $L=\{L_{1},\dots,L_{N}\}$.
2:A collaborative knowledge graph $\mathrm{KG}_{\text{collab}}$.
3:Initialize: $\mathrm{KG}_{\text{explicit}}\leftarrow\emptyset$, $\mathrm{KG}_{\text{tacit}}\leftarrow\emptyset$
4:
[…]
6:for $i\leftarrow 1$ to NN do
7:  for all $(\text{instruction},\text{response})\in D_{i}$ do
8:   $\mathrm{triples}\leftarrow\text{ExtractTriples}(\text{instruction},\text{response})$
9:   $\text{LabelNodesInTriples}(\mathrm{triples},L_{i})$
10:   $\mathrm{KG}_{\text{explicit}}\leftarrow\mathrm{KG}_{\text{explicit}}\cup\mathrm{triples}$
11:  end for
12:end for
13:
[…]
15:for $i\leftarrow 1$ to NN do
16:  for all $(\text{instruction},\text{response})\in D_{i}$ do
17:   $\mathrm{cf\_instruction}\leftarrow\text{GenerateCounterfactual}(\text{instruction})$
18:   repeat
19:   $\mathrm{cf\_response}\leftarrow A_{i}.\text{query}(\mathrm{cf\_instruction})$
20:   $(\text{evaluation},\text{feedback})\leftarrow E.\text{evaluate}(\mathrm{cf\_response})$
21:   if $\text{evaluation}\neq\text{"reasonable"}$ then
22:     $\mathrm{cf\_instruction}.\text{append}(\text{feedback})$
23:   end if
24:   until $\text{evaluation}=\text{"reasonable"}$
25:   $\mathrm{tacit\_triples}\leftarrow\text{ExtractTriples}(\mathrm{cf\_instruction},\mathrm{cf\_response})$
26:   $\text{LabelNodesInTriples}(\mathrm{tacit\_triples},L_{i})$
27:   $\mathrm{KG}_{\text{tacit}}\leftarrow\mathrm{KG}_{\text{tacit}}\cup\mathrm{tacit\_triples}$
28:  end for
29:end for
30:
[…]
33:return $\mathrm{KG}_{\text{collab}}$
Algorithm 2 Causal Chain Retrieval and Construction
1:User query $\mathrm{query}$; Stage-labeled knowledge graph $\mathrm{KG}_{\text{collab}}$; Text encoder $\mathrm{Encoder}$; Seed nodes count $\mathrm{top\_n}$; Chain length limit $\mathrm{max\_depth}$.
2:A textual causal chain $\mathrm{causal\_chain\_text}$.
3:
[…]
11:
[…]
14:for each $\mathrm{start\_node}$ in $\mathrm{seed\_nodes}$ do
15:  $\triangleright$ Search paths crossing stages via bridge nodes with multiple labels
16:  $\mathrm{paths}\leftarrow\mathrm{KG}_{\text{collab}}.\text{FindCrossStagePaths}(\mathrm{start\_node},\mathrm{min\_depth}\leftarrow 2,\mathrm{max\_depth})$
17:  $\mathrm{all\_paths}\leftarrow\mathrm{all\_paths}\cup\mathrm{paths}$
18:end for
19:
[…]
22:return $\mathrm{causal\_chain\_text}$
Algorithm 3 Prompts Tree Construction
1:Workflow stages $S=\{S_{1},\dots,S_{n}\}$; Domain agents $A=\{A_{1},\dots,A_{n}\}$; Seed Q-A pair $\mathrm{InitialSeedQA}$; Best prompts count $m$.
2:A prompt tree $\mathrm{PromptsTree}$.
3:Initialize:
[…]
7:while $\mathrm{queue}\neq\emptyset$ do
8:  $\mathrm{parentNode},\mathrm{currentAnswer},\mathrm{currentStage}\leftarrow\mathrm{queue}.\text{pop\_front}()$
9:  $\triangleright$ Phase 1: Distill Prompts
10:  $\mathrm{distilledPrompts}\leftarrow A_{\mathrm{currentStage}}.\text{DistillPrompts}(\mathrm{currentAnswer})$
11:  $\triangleright$ Phase 2: Select Best Prompts
12:  $\mathrm{bestPrompts}\leftarrow A_{\mathrm{currentStage}}.\text{SelfEvaluate}(\mathrm{distilledPrompts},m)$
13:  $\triangleright$ Phase 3: Grow Tree
14:  if $\mathrm{currentStage}\neq S_{n}$ then $\triangleright$ Proceed if not the last stage
15:   $\mathrm{nextStage}\leftarrow S_{\text{index}(\mathrm{currentStage})+1}$
16:   for each $\mathrm{promptText}$ in $\mathrm{bestPrompts}$ do
17:     $\mathrm{childNode}\leftarrow\text{CreateNode}(\mathrm{promptText})$
18:     $\mathrm{parentNode}.\text{add\_child}(\mathrm{childNode})$
19:     $\mathrm{newQuestion}\leftarrow\text{GenerateQuestionForNextStage}(\mathrm{promptText},\mathrm{nextStage})$
20:     $\mathrm{newAnswer}\leftarrow A_{\mathrm{nextStage}}.\text{query}(\mathrm{newQuestion})$
21:     $\mathrm{queue}.\text{push\_back}((\mathrm{childNode},\mathrm{newAnswer},\mathrm{nextStage}))$
22:   end for
23:  end if
24:end while
25:return $\mathrm{PromptsTree}$
As shown in the Figure 9, after applying Cochain, the agent anticipates that carbon fiber requires complex molding processes, completely different from traditional metal stamping, and that qualified suppliers are scarce, thus requiring consideration of supply chain stability and high procurement costs significantly impacting vehicle pricing and profit margins. The prompts tree anticipates existing production processes, guiding the agent to consider compatibility constraints between new equipment and existing stamping-welding lines. Based on this, Cochain generates superior answers compared to single agents.
Figure 8: Examples of implementing a core component in the Cochain framework.
Figure 9: Examples of single agent and complete workflow of Cochain
Appendix B Proof: Tacit Knowledge & Counterfactual Output Model
Given the inherent difficulty in directly extracting hidden layer knowledge from agents operating within complex systems, our approach introduces counterfactual reasoning as a means to probe and subsequently infer tacit knowledge. This section elucidates the probabilistic formulations developed to model these inferred internal aspects of each agent. It details the conceptualization of tacit knowledge, the generation of counterfactual outputs by the agent, and the overall generative process used to understand these unobservable agent workings.
The foundation of our approach to modeling tacit knowledge within each agent lies in its conditional generation based on latent tacit variables. We define $h_{ij}$ as the tacit knowledge specific to a sample $(X_{ij},Y_{ij})$ processed by a particular agent. Within an agent, $h_{ij}$ represents an internal state or representation that is not directly accessible. While $h_{ij}$ is an abstract construct for modeling these unobservable aspects, for illustrative purposes, one might conceptualize it as akin to intermediate computational states, learned feature representations, or internal configurations within the Agent that influence its output generation. Concurrently, $\theta_{ij}$ is defined as a latent tacit variable associated with the same sample and agent. This variable $\theta_{ij}$ is also an abstract construct, encapsulating underlying factors hypothesized to influence $h_{ij}$, such as specific input characteristics processed by the Agent, its operational modes, or conditioning contexts it operates under. The introduction of $\theta_{ij}$ and $h_{ij}$ is an attempt to create a probabilistic model for the Agent’s internal, unobservable precursors to decision-making or generation.
A core assumption of our model is that this tacit knowledge $h_{ij}$ is not deterministically derived but is rather a stochastic realization conditioned on $\theta_{ij}$. This is particularly relevant for complex Agents, where output generation can be influenced by internal stochastic mechanisms or intricate decision processes. This probabilistic relationship is formally expressed as:
$h_{ij}\mid\theta_{ij}\sim P(h_{ij}\mid\theta_{ij})$ (8)
This notation signifies that $h_{ij}$ is a random variable whose probability distribution $P(h_{ij}\mid\theta_{ij})$ is conditional upon the specific value or state of $\theta_{ij}$. The nature of this distribution would be specific to the agent and reflects the inherent complexities and uncertainties in how its internal states are formed. For instance, $\theta_{ij}$ might represent an abstract control signal or a high-level interpretation of input patterns by the agent, and $P(h_{ij}\mid\theta_{ij})$ would describe the distribution of possible internal representations resulting from it. This formulation acknowledges that even with a defined influencing factor $\theta_{ij}$, the precise tacit knowledge $h_{ij}$ (the agent’s internal processing focus or state) can vary. The characteristics of $P(h_{ij}\mid\theta_{ij})$ can be explored and potentially learned through systematic counterfactual probing of the agent.
Following the generation of tacit knowledge, we model the generation of the counterfactual output $\tilde{Y}_{ij}$ by the agent. This output is conditioned not only on the counterfactual input $\tilde{X}_{ij}$ but also critically on the realized tacit knowledge $h_{ij}$ derived from Equation 8. We define $\tilde{Y}_{ij}$ as the Agent’s generated output when presented with the counterfactual input $\tilde{X}_{ij}$, and $h_{ij}$ acts as the mediating internal state through which $\tilde{X}_{ij}$ is processed to produce $\tilde{Y}_{ij}$.
The generation of $\tilde{Y}_{ij}$ by such agents is often an inherently probabilistic process. Even with a given counterfactual input $\tilde{X}_{ij}$ and a specific internal tacit knowledge state $h_{ij}$, the agent may not produce a single, deterministic output. This inherent stochasticity is captured by the following conditional probability distribution, whose parameters might also be inferred from counterfactual observations:
$\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij}\sim P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ (9)
This expression states that $\tilde{Y}_{ij}$ (the counterfactual output) is a random variable whose distribution $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ depends on both the specific tacit knowledge state $h_{ij}$ and the counterfactual input $\tilde{X}_{ij}$. The tacit knowledge $h_{ij}$ here serves as a crucial intermediary. For agents producing sequential outputs, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$ might be represented as an autoregressive factorization, $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})=\prod_{t}P(\tilde{y}_{ij,t}\mid\tilde{y}_{ij,<t},h_{ij},\tilde{X}_{ij})$.
To determine the overall probability of observing a counterfactual output $\tilde{Y}_{ij}$ from an agent given only the counterfactual input $\tilde{X}_{ij}$—a common scenario when internal states are not directly observable—we must account for the entire hypothesized generative pathway. This involves marginalizing out the unobserved latent tacit variable $\theta_{ij}$ and, by extension, the intermediate tacit knowledge $h_{ij}$. The final generation process is thus expressed as:
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$ (10)
This equation can be derived by considering the dependencies established in Equations 8 and 9.
The derivation proceeds as follows: First, we apply the law of total probability to marginalize over the latent tacit variable $\theta_{ij}$. The probability of $\tilde{Y}_{ij}$ given $\tilde{X}_{ij}$ can be written as an integral over all possible values of $\theta_{ij}$ in its domain $\Theta$:
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$
Using the definition of conditional probability, $P(A,B\mid C)=P(A\mid B,C)P(B\mid C)$, we can rewrite the integrand $P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})$ as:
$P(\tilde{Y}_{ij},\theta_{ij}\mid\tilde{X}_{ij})=P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})$
Here, $P(\theta_{ij}\mid\tilde{X}_{ij})$ represents the prior probability distribution of the tacit variable $\theta_{ij}$ given the counterfactual input $\tilde{X}_{ij}$. This term reflects how a given counterfactual input might stochastically lead to different internal configurations or “operational modes” $\theta_{ij}$ within the Agent that are not directly observable. The nature of this prior can also be investigated through counterfactual analysis. Substituting this back, we get:
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$ (11)
The term $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$ is the probability of generating $\tilde{Y}_{ij}$ given both $\theta_{ij}$ and $\tilde{X}_{ij}$. Our model posits a specific structure for this, reflecting a Markov chain-like dependency: $\theta_{ij}\rightarrow h_{ij}\rightarrow\tilde{Y}_{ij}$ (all conditioned on $\tilde{X}_{ij}$ where appropriate). This structure implies two key conditional independence assumptions:
  1. 1.
    Given $\theta_{ij}$, $h_{ij}$ is independent of $\tilde{X}_{ij}$ if all influence of $\tilde{X}_{ij}$ on $h_{ij}$ is mediated through $\theta_{ij}$. More commonly, $P(h_{ij}\mid\theta_{ij})$ is defined as the direct influence of $\theta_{ij}$ on $h_{ij}$.
  2. 2.
    Given $h_{ij}$ and $\tilde{X}_{ij}$, $\tilde{Y}_{ij}$ is independent of $\theta_{ij}$. That is, $P(\tilde{Y}_{ij}\mid h_{ij},\theta_{ij},\tilde{X}_{ij})=P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})$. The tacit knowledge $h_{ij}$ fully mediates the influence of $\theta_{ij}$ on $\tilde{Y}_{ij}$.
Under these assumptions, we can decompose $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$:
$P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})=\int P(\tilde{Y}_{ij},h_{ij}\mid\theta_{ij},\tilde{X}_{ij})dh_{ij}$
$\int P(\tilde{Y}_{ij}\mid h_{ij},\theta_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij},\tilde{X}_{ij})dh_{ij}$
Applying the conditional independence assumptions:
$\int P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})dh_{ij}$
The form in Equation 10 implies that for a given $\theta_{ij}$ in the outer integral, the relevant $h_{ij}$ is the one generated from that $\theta_{ij}$, effectively collapsing the inner integral. Thus, $P(\tilde{Y}_{ij}\mid\theta_{ij},\tilde{X}_{ij})$ is represented by the product $P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})$, where $h_{ij}$ is the specific realization tied to $\theta_{ij}$ within the integral. Substituting this construction into Equation 11 directly yields Equation 10:
$P(\tilde{Y}_{ij}\mid\tilde{X}_{ij})=\int_{\Theta}P(\tilde{Y}_{ij}\mid h_{ij},\tilde{X}_{ij})P(h_{ij}\mid\theta_{ij})P(\theta_{ij}\mid\tilde{X}_{ij})d\theta_{ij}$
This integral sums the probabilities of $\tilde{Y}_{ij}$ occurring via all possible latent tacit variables $\theta_{ij}$. Each path’s contribution is weighted by the prior probability of $\theta_{ij}$, the probability of the corresponding tacit knowledge $h_{ij}$ arising from that $\theta_{ij}$, and finally the probability of the output $\tilde{Y}_{ij}$ given that $h_{ij}$ and the input $\tilde{X}_{ij}$.
Appendix C Implementation Details
Definition of Business Workflow: It is a complex process with many interconnected stages. These stages are specialized. The key feature is that these stages strongly depend on each other. A decision in one stage requires more than just expert knowledge of that area. It must also anticipate and use knowledge and constraints from the other stages.
Table 6 presents a comprehensive comparison of Cochain against other baseline models. Cochain demonstrates superior agent collaboration, characterized by faster inference speeds and lower hallucination rates. More critically, Cochain exhibits decomposability, ensuring that the collaboration process is resilient and does not terminate due to failures in individual components. It is important to note that for CoA, source documents are not included due to the differing task orientation.
The LLMs referenced in Figure 5 (Main Paper) are fine-tuned using QLoRA Dettmers et al. [2023], configured with a LoRA rank of 8 and trained for 3 epochs, on two H100 GPUs, with models quantized to 4-bit precision via bitsandbytes. Other open-source models are fine-tuned for 10 epochs using LoRA Hu et al. [2022]. The inference times reported in Figure 10 and Table 7 are benchmarked by deploying the open-source models on a single A40 GPU to mitigate network variability.
Table 6: Comparison of different multi-agent collaboration methods
Baseline Collaboration Method Decomposable Reasoning Speed Hallucination Rate Agent
PMC Zhang et al. [2025a] Hierarchical Planning Low Middle Multiple
MedAgents Tang et al. [2024] Discussion & Voting Low Middle Multiple
Debate(short) Du et al. [2024] Debate Low Middle Multiple
Debate(long) Du et al. [2024] Debate Low Middle Multiple
CoA Zhang et al. [2024b] Chain Collaboration Low Middle Multiple
Original - High High Single
Cochain Knowledge Fusion High Low Multiple
The appendix provides a more extensive report on experimental results across additional metrics. These include BERTScore-P, BERTScore-R, BLEU-4 Papineni et al. [2002] for measuring text alignment, METEOR Banerjee and Lavie [2005] which considers semantically equivalent phrases, ROUGE-2 for assessing content breadth, and ROUGE-L Lin [2004] which evaluates deep semantic alignment, particularly for long texts.
Appendix D More Experiments and Analysis
D.1 More Efficiency and Cost Analysis
Table 7 presents a comparative analysis of Cochain against other multi-agent systems, focusing on key operational metrics: Inference Time, Token Throughput, and Cost. A noteworthy observation on the Automotive and E-commerce datasets is that Cochain’s average input token count is approximately one-twentieth of that utilized by PMC and CoA. This disparity suggests that these alternative systems expend a considerable volume of tokens on inter-agent communication. In contrast, Cochain innovatively transforms such communication into structured knowledge graphs and prompts tree, a mechanism that effectively curtails operational costs. The overall results compellingly demonstrate that Cochain achieves substantial reductions in inference time, token throughput, and aggregate cost when benchmarked against other multi-agent approaches. Remarkably, the efficiency metrics for Cochain closely approximate those observed for single-agent systems, and its average cost is even lower, underscoring a significant advantage of our proposed framework in terms of computational resource utilization and operational expenditure. Table 8 presents the retrieval contribution of each component during the inference process. The prompt tree retrieval employs a depth-first search (DFS) strategy. In terms of time complexity, this is equivalent to a simple path traversal with a complexity of $O(D)$, where $D$ denotes the depth of the business process. Figure 10 shows that as reasoning rounds increase, the time gap between thinking and non-thinking models grows dramatically for PMC but remains moderate for Cochain, demonstrating Cochain’s superior efficiency as complexity increases.
Figure 10: ExampReasoning time differential.
Table 7: Comparative analysis of inference time, token throughput, and cost. Inference time is measured using a locally deployed Qwen2-7B model to obviate network latency associated with API calls. Conversely, token throughput and cost metrics are derived using Claude-3.5-Haiku.
Baselines Inference Time (s) Avg.Input Avg.Output Avg.Cost (\$)
Automotive
PMC Zhang et al. [2025a] 241.56 11637.94 6644.22 0.0359
MedAgents Tang et al. [2024] 196.53 3901.03 3483.24 0.0170
Debate(short) Du et al. [2024] 187.16 3938.35 4809.27 0.0223
Debate(long) Du et al. [2024] 162.57 4357.87 5278.53 0.0246
CoA Zhang et al. [2024b] 227.21 11428.06 5446.62 0.0309
Original 37.22 93.49 1041.38 0.0042
Cochain 38.77 501.62 541.74 0.0026
Pharmaceutical
PMC Zhang et al. [2025a] 198.49 3796.07 2613.73 0.0135
MedAgents Tang et al. [2024] 169.53 3504.85 3504.85 0.0168
Debate(short) Du et al. [2024] 548.66 3372.56 3772.79 0.0178
Debate(long) Du et al. [2024] 574.95 3810.87 4268.74 0.0201
CoA Zhang et al. [2024b] 253.18 5493.94 3450.45 0.0182
Original 60.96 77.28 1091.32 0.0044
Cochain 44.70 533.88 830.34 0.0037
E-commerce
PMC Zhang et al. [2025a] 256.84 11319.74 6452.89 0.0349
MedAgents Tang et al. [2024] 181.61 3816.26 3445.34 0.0169
Debate(short) Du et al. [2024] 318.51 4414.23 5660.94 0.0261
Debate(long) Du et al. [2024] 293.42 5438.52 6911.85 0.0321
CoA Zhang et al. [2024b] 195.49 11446.05 5353.24 0.0306
Original 35.83 60.90 1080.92 0.0044
Cochain 37.10 561.80 620.15 0.0029
Table 8: Latency Breakdown Analysis of Cochain’s Inference Components
Inference Component Average Time (s) Percentage of Total Time (%)
Knowledge Graph Retrieval 0.6200 s 1.57 %
Prompts Tree Retrieval 0.0022 s <0.01 %
LLM Model Inference 38.7700 s 98.42 %
D.2 More Experiments on Single Agent
Table 9 and Table 10 demonstrate that integrating Cochain markedly enhances single-agent performance across diverse backbone models during two distinct stages of a business workflow. This enhancement is consistently reflected across all evaluated metrics, irrespective of whether the LLMs are in their raw, pre-trained state or have undergone fine-tuning. Notably, a synergistic effect between fine-tuning and the application of Cochain typically yields optimal performance, underscoring Cochain’s robust capability to leverage specialized model knowledge. The broad-spectrum improvements observed across various model metrics highlight the effectiveness and versatility of Cochain in mitigating prior issues of insufficient collaboration.
Table 9: The performance of our proposed method on single-agent. BERTScore F1 is abbreviated as BS-F, BERTScore Precision as BS-P, and BERTScore Recall as BS-R. We highlight the best and second-best results.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Raw LLM 62.25 64.41 60.54 6.04 8.28 17.52 7.78 15.64
+ Cochain 65.70 68.25 63.50 10.99 13.84 25.50 10.44 16.46
Finetuned LLM 70.68 70.71 70.67 14.31 18.79 35.03 11.97 21.49
+ Cochain 71.46 71.35 71.59 16.50 20.96 38.69 13.79 21.58
Qwen2-7B Raw LLM 71.76 71.79 71.75 17.23 22.30 41.13 12.88 24.00
+ Cochain 75.24 74.90 75.58 22.32 26.51 52.00 19.53 27.94
Finetuned LLM 71.72 71.75 71.70 16.90 21.64 40.77 13.35 24.07
+ Cochain 75.39 75.14 75.65 24.17 28.17 51.22 21.22 28.87
Llama3-8B Raw LLM 68.93 68.38 69.59 7.47 9.14 24.48 10.99 19.28
+ Cochain 73.75 73.36 74.17 19.47 23.52 48.43 17.57 25.10
Finetuned LLM 70.89 70.88 70.91 14.90 19.05 39.83 12.59 22.90
+ Cochain 74.56 74.13 75.01 22.12 26.15 50.59 19.41 27.19
Llama2-13B Raw LLM 60.79 63.16 59.00 5.73 7.96 16.90 7.52 15.43
+ Cochain 67.47 70.20 65.15 13.42 16.88 29.97 13.65 18.57
Finetuned LLM 70.88 70.88 70.90 14.57 19.14 35.87 12.19 21.85
+ Cochain 71.80 72.08 71.58 18.16 22.29 40.33 14.96 23.18
Qwen2.5-14B Raw LLM 71.86 72.03 71.70 9.48 12.16 40.71 11.85 17.86
+ Cochain 75.62 75.56 75.69 12.39 14.71 47.95 17.66 19.55
Finetuned LLM 71.50 71.48 71.53 9.25 11.75 40.16 12.01 17.67
+ Cochain 75.18 75.00 75.38 13.11 15.08 48.35 19.70 20.30
Table 10: The performance of our proposed method. The results of this table are in a different stage, as compared to the stage presented in Table 9. We highlight the best and second-best results.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Raw LLM 61.76 64.72 59.44 6.36 8.69 18.25 8.66 16.17
+ Cochain 63.57 66.31 61.23 9.49 12.37 23.12 8.99 17.46
Finetuned LLM 69.88 69.78 70.00 12.02 17.02 32.59 9.26 20.69
+ Cochain 70.47 70.43 70.55 13.79 18.58 35.44 10.91 20.26
Qwen2-7B Raw LLM 70.34 71.12 70.60 13.74 19.88 35.29 9.47 22.01
+ Cochain 73.48 72.37 74.63 15.77 19.21 50.38 16.25 21.99
Finetuned LLM 70.28 70.08 70.50 13.61 18.95 35.07 9.96 22.38
+ Cochain 74.51 74.30 74.74 21.99 25.76 50.49 20.71 26.00
Table 11: Quantifying insufficient collaboration.
Backbone Method BS-F GLEU ROUGE-L
Qwen2-7B Single Agent 71.72 21.64 24.07
+ Manual Prompts 72.54 ($\uparrow$) 21.83 ($\uparrow$) 24.34 ($\uparrow$)
+ Cochain 75.39 ($\Uparrow$) 28.17 ($\Uparrow$) 28.87 ($\Uparrow$)
DeepSeek-R1-7B Single Agent 71.35 19.55 22.71
+ Manual Prompts 71.83 ($\uparrow$) 19.90 ($\uparrow$) 23.38 ($\uparrow$)
+ Cochain 73.19 ($\Uparrow$) 21.71 ($\Uparrow$) 24.01 ($\Uparrow$)
As shown in Table 11, to better quantify the impact of insufficient collaboration, we compare the baseline single agent to one provided with minimal communication, simulated by manually injecting five prompts containing cross-domain constraints. The results of this table clearly show: First, insufficient collaboration is a real problem, as the agent with injected communication showed a slight performance improvement over the original agent. This directly proves that adding cross-domain communication improves outcomes, thereby quantifying the performance loss from insufficient collaboration. Second, Cochain’s intelligent collaboration is far superior to simple communication, as the performance improved substantially after applying Cochain. This demonstrates that Cochain’s value is not just about adding information. Rather, it excels at intelligently filtering, organizing, and structuring collaborative knowledge. Its effectiveness far surpasses that of unstructured, manually added information and helps to avoid the risk of excessive collaboration.
D.3 More Experiments on Specialized Agent Collaboration
Figure 6 (Main Paper) presents the results from a distinct experimental configuration designed to further investigate the coordination capabilities of Cochain. Whereas Table 9 assessed a multi-domain model (a single Qwen2-7B model fine-tuned on all domains) and specialized models (different models each fine-tuned for specific domains), Table 12 exclusively focuses on a setup utilizing multiple agents of the same backbone, each independently fine-tuned for a specific domain. The findings indicate that Cochain effectively facilitates collaboration among these identically architected yet uniquely specialized agents. Performance enhancements are observed both within their respective specialized domains and in the quality of the comprehensive, integrated solutions. This underscores Cochain’s distinct advantage in synergizing specialized knowledge from a team of uniquely trained agents that share a common architectural foundation.
Table 12: Results for one model at particular stages. This means that we use Qwen2-7B for four fine-tunings, and the four fine-tuned models are experimented with at their respective stages. We highlight the fine-tuned results, Cochain application, and improvements()($\uparrow$).
Backbone Stage Method BS-F BS-P BS-R BLEU-4 GLEU ROUGE-L
Qwen2-7B $S_{1}$ Finetuned 71.72 71.75 71.70 16.90 21.64 24.07
+ Cochain 75.39 75.14 75.65 24.17 28.17 28.87
$\uparrow$ +3.67 +3.39 +3.95 +7.27 +6.53 +4.80
$S_{2}$ Finetuned 70.24 70.47 70.16 13.19 18.29 20.95
+ Cochain 74.23 74.01 74.46 20.05 23.57 24.71
$\uparrow$ +3.99 +3.54 +4.30 +6.86 +5.28 +3.76
$S_{3}$ Finetuned 70.28 70.08 70.50 13.61 18.95 22.38
+ Cochain 74.51 74.30 74.74 21.99 25.76 26.00
$\uparrow$ +4.23 +4.22 +4.24 +8.38 +6.81 +3.62
$S_{4}$ Finetuned 69.40 69.49 69.32 13.51 18.26 21.82
+ Cochain 75.06 74.99 75.15 23.95 27.70 27.49
$\uparrow$ +5.66 +5.50 +5.83 +10.44 +9.44 +5.67
Score 6.4 9.5 +3.1
D.4 More Results from Ablation Studies
On the automotive dataset, we report additional ablation study results focusing on the impact of different backbone models. As indicated in Table 13, “w/o Prompts Tree” markedly impaired performance. Specifically, when applied to the Llama-8B backbone, this ablation led to a 1.97% reduction in the BS-F score, with notable degradations also observed across other evaluation metrics. These findings are consistent with the experimental results presented in Table 5 (Main Paper), further underscoring the critical role of the Prompts Tree component in our framework.
Table 13: More results of the ablation study. $\mathcal{CKG}$ represents the collaborative knowledge graph. We highlight the most and second-most efficient modules.
Backbone Method BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
Llama2-7B Cochain 71.46 71.35 71.59 16.50 20.96 38.69 13.79 21.58
w/o $\mathcal{CKG}$ 71.26 71.14 71.41 16.06 20.53 38.21 13.55 21.50
w/o Causal Chain 71.26 71.13 71.44 15.63 19.59 40.32 14.33 21.16
w/o Prompts Tree 71.06 70.92 71.20 15.20 19.61 36.66 13.07 21.38
Qwen2-7B Cochain 75.39 75.14 75.65 24.17 28.17 51.22 21.22 28.87
w/o $\mathcal{CKG}$ 75.21 74.91 75.53 23.56 27.81 49.76 20.15 28.74
w/o Causal Chain 75.20 74.92 75.50 23.68 27.55 51.00 21.14 28.42
w/o Prompts Tree 73.08 73.20 72.98 18.40 22.45 45.43 16.06 24.56
Llama3-8B Cochain 74.56 74.13 75.01 22.12 26.15 50.59 19.41 27.19
w/o $\mathcal{CKG}$ 74.27 73.80 74.75 21.45 25.46 49.99 19.09 27.13
w/o Causal Chain 73.63 72.79 74.50 17.34 20.70 49.15 18.28 23.78
w/o Prompts Tree 72.59 72.45 72.75 17.01 21.02 45.00 15.18 23.83
Qwen2.5-14B Cochain 75.18 75.00 75.38 13.11 15.08 48.35 19.70 20.30
w/o $\mathcal{CKG}$ 74.83 74.41 75.25 12.49 14.55 47.32 18.39 19.76
w/o Causal Chain 74.71 74.66 74.79 12.28 14.17 47.65 19.30 19.44
w/o Prompts Tree 73.01 73.01 73.02 9.98 12.25 42.72 14.15 18.00
D.5 Integration of Cochain with Baselines
We further investigate the integration of Cochain with existing baseline methods, specifically PMC and MedAgents, through experiments conducted on a pharmaceutical dataset using Claude-3.5-Haiku. As reported in Table 14, the combination of Cochain with PMC results in improvements of 2.62% and 7.55% in the BS-F and BLEU-4 scores, respectively. This outcome further substantiates the efficacy of Cochain in mitigating issues associated with excessive collaboration.
Table 14: The improvement of Cochain over other baselines on pharmaceutical datasets.
Baseline BS-F BS-P BS-R BLEU-4 GLEU METEOR ROUGE-2 ROUGE-L
PMC Zhang et al. [2025a] 66.51 66.06 66.97 5.17 12.33 20.31 5.48 13.56
PMC Zhang et al. [2025a] + Cochain 69.13 68.49 69.80 12.72 18.48 27.01 6.93 15.82
MedAgents Tang et al. [2024] 66.23 66.02 66.48 7.17 14.02 22.59 5.43 14.32
MedAgents Tang et al. [2024] + Cochain 67.45 66.91 68.00 11.65 17.80 27.06 6.16 15.74
D.6 Evaluation of Collaborative Knowledge Graph
To evaluate the reliability of our final knowledge graph, we randomly sampled 2,000 triplets from it, consisting of 1,000 from explicit knowledge extraction and 1,000 from our counterfactual method. We used three core metrics for evaluation: Factual Correctness (Is the knowledge factually correct in the real world?),Task Relevance (How helpful is the knowledge for solving business workflow problems?), and Knowledge Depth (Is the knowledge common sense, or does it require professional insight?) We invited two domain experts for a blind review. We also used Gemini-2.5-Pro as a judge for an objective evaluation. The average scores (out of 5) are presented in the Table 15.
Table 15: Evaluation of Collaborative Knowledge Graphs by Human Experts and LLMs.
Evaluation Metric Human Expert Score (Avg. / 5.0) Gemini-2.5-Pro Score (Avg. / 5.0)
Factual Correctness 4.99 5.00
Task Relevance 4.76 4.81
Knowledge Depth 4.31 4.54
D.7 Excessive collaboration is More Serious than Insufficient collaboration
Business workflow tasks are particularly susceptible to the “excessive collaboration” phenomenon. When multiple agents collaborate, this manifests as responses deviating from core issues and reducing answer quality. Comparative analysis of Table 1 (Main Paper), Figure 6 (Main Paper), and additional single-agent experimental results (Supplementary Material D.2) demonstrates that, despite consuming significantly more computational resources, excessive collaboration performs substantially worse than insufficient collaboration. excessive collaboration’s challenging nature is its covertness—systems maintain high activity and apparent collaboration, making the problem hard to detect and rectify promptly. While task decomposition, such as approaches like PMC, and summarization, such as strategies like CoA, are considered effective methods for focusing on core issues Wu et al. [2024]; Chen et al. [2024], Cochain’s collaboration mechanism exhibits superior performance in controlling excessive collaboration, enabling a more precise focus on critical task requirements.
D.8 Robustness to Incomplete Retrieval
To validate the system’s robustness against potential keyword retrieval failures, we conducted a stress test. We designed 50 structured queries in the format of “How to use [Core Concept A] to solve problems related to [Target Concept B]”. In the test condition, we deliberately masked all keywords corresponding to [Target Concept B] to simulate a first-stage retrieval failure. For instance, in the query “How can ergonomics be used to reduce traffic accident injuries?”, the keywords “traffic” and “accident” were blocked.
While end-to-end semantic retrieval could theoretically solve keyword-mismatch issues, it becomes computationally prohibitive for real-time interaction on massive knowledge graphs. We therefore adopt the Retrieval-Ranking two-stage paradigm, a proven industry standard Covington et al. [2016]. This approach combines the advantages of both methods: keyword-based retrieval in the first stage ensures precise matching of technical terms and proper nouns, avoiding omissions from semantic over-generalization. In the second stage, semantic ranking captures the query’s overall intent, ensuring system generalization and robustness.
Table 16: Performance comparison under simulated retrieval failure. Masking the target concept’s keywords results in a negligible performance drop, demonstrating the system’s robustness.
Test Condition BS-F GLEU ROUGE-L
Full Query 74.16 22.08 25.86
Target Concept Masked 73.87 21.42 24.57
Performance Drop (%) -0.29% -0.66% -1.29%
The results in Table 16 confirm the effectiveness of this design. Even when forcing a first-stage failure, the BS-F score experienced a minimal drop of only 0.29%, from 74.16 to 73.87, with similarly small decreases in GLEU (-0.66%) and ROUGE-L (-1.29%) scores. This resilience is attributed to two factors. First, as justified above, the semantic ranking stage corrects for an imperfect initial candidate set. More critically, the subsequent graph traversal and causal chain stages provide a powerful secondary path to completeness. Even if “Traffic Accident” is missed during retrieval, it can be discovered through pre-existing knowledge paths represented as triples (e.g., (Injuries, is_a_result_of, Traffic Accident)), ensuring the final context remains robust and comprehensive.
Appendix E Case Study
E.1 Interpretability Analysis
We investigate the interpretability characteristics of the model after the application of Cochain through a case study. As shown in Table 17, we employ the same color to annotate logically related content visually. The findings reveal that, following the application of Cochain, the model is capable of performing multi-dimensional reasoning based on knowledge graphs, causal inference chains, and contextual cues when addressing user needs, thereby generating outputs that are traceable, reliable, and interpretable. The application of the Cochain method has significantly enhanced the transparency and explainability of the model’s reasoning process, providing empirical evidence for understanding the model’s collaborative decision-making mechanisms.
E.2 Comparative Analysis of Model Outputs
To understand how different baselines handle complex instructions, as shown in Figure 11, we qualitatively analyzed their outputs against a reference answer, focusing on their integration of User Experience(UX) design within the automotive business workflow. Additionally, we present case studies for the pharmaceutical business workflow and the e-commerce business workflow, as shown in Figure 12 and Figure 13.
Cochain: Clear Structure, Focused on UX Integration Process.
Cochain’s response was notably well-structured, using a seven-point list for key UX integration stages, enhancing readability and providing a clear action framework. It comprehensively covered the UX lifecycle—from user research and design principles to interdisciplinary collaboration, technological innovation, continuous iteration, and performance-cost/quality assurance. Its explicit “collaborative perspective of the automotive business workflow” aligned with the reference answer’s holistic view, effectively capturing its spirit by offering a structured, UX-centric methodology.
Comparison of Cochain with Other Model Responses.
Cochain’s primary strength is its sustained focus on the UX design integration process and principles, closely matching the prompt’s core intent and the reference answer. In contrast, PMC and CoA lean towards high-level strategic planning (e.g., PLM platforms, KPIs), emphasizing system management rather than the deep UX integration across workflow stages highlighted by the reference answer. Debate showcases significant depth in UX concepts and cutting-edge technologies. However, its primary focus is on HMI innovation details, rather than systematically integrating UX across broader business workflow processes. Cochain can better balance innovation with this workflow integration. MedAgents share some structural similarities with Cochain, but Cochain is more direct in articulating the end-to-end UX lifecycle, especially “Continuous Iteration and Optimization,” and its business workflow perspective. While other models excel in specific areas (e.g., Debate’s UX tech depth; PMC and CoA’s strategic scope), Cochain demonstrates the closest overall alignment with the reference answer’s advocated UX-integrated process throughout the business workflow and its pragmatic perspective, thus performing excellently in this qualitative assessment.
Table 17: Case study of interpretability. The three distinct modules, namely Knowledge, Causal Chain, and Prompts Tree, are each assigned one of the three different color schemes: red, yellow, and blue. The colors within the modules correspond to those in the output, and logically related content is annotated using the same color scheme.
User need How can cost optimization be achieved in automotive manufacturing?
Knowledge Cost optimization points include technological innovation, quality control, and environmental and energy management. Design optimization, material selection, and supply chain network construction strategies are measures for cost optimization…
Cause chain Stamping is primarily used for component forming and can be integrated with automated and intelligent production lines, intelligent monitoring and optimization systems, as well as smart mold maintenance
Prompts tree Prompt 1: User Need: Cost optimization $\rightarrow$ Design Method: Accelerating prototype iteration $\rightarrow$ Supply Chain Method: Selection of lightweight materials $\rightarrow$ Production Method: Research and analysis of lightweight material performance and cost $\rightarrow$ Quality Inspection Method: ensuring materials meet performance and cost requirements.
Prompt 2: User Need: Cost optimization $\rightarrow$ Design Method: Accelerating prototype iteration $\rightarrow$ Supply Chain Method: Selection of lightweight materials $\rightarrow$ Production Method: Design optimization to enhance material utilization $\rightarrow$ Quality Inspection Method: Testing the strength and stability of the optimized structure.
Output In automotive design, it is essential to consider prototype optimization and innovation in design methodologies. For instance, to address customer-specific vehicle customization, cost control during the design phase can be achieved by accelerating prototype iterations and utilizing lightweight materials. Additionally, the properties and costs of lightweight materials should be taken into account to optimize structural design and enhance material utilization. In supply chain management, cost optimization can also be achieved. This includes strategies for building the supply chain network and optimizing logistics. By consolidating procurement and leveraging economies of scale, procurement costs can be reduced. Collaborative Planning, Forecasting, and Replenishment (CPFR) methods can further optimize inventory costs. In the manufacturing stage, cost optimization can be achieved through the introduction and implementation of automated and intelligent production lines. For example, the use of automated production lines, coupled with intelligent monitoring and optimization systems, can reduce redundant tasks and increase production efficiency. Furthermore, for specific manufacturing processes, such as stamping, material performance costs can be analyzed, and intelligent molds and automatic maintenance systems can be developed to optimize costs. Finally, in the quality inspection stage, ensuring that all materials and products meet the predetermined performance and cost requirements is crucial. Additionally, improving production processes to ensure product processing quality is an important aspect of quality inspection.
Figure 11: Case study on the Automotive dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Figure 12: Case study on the Pharmaceutical dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Figure 13: Case study on the E-commerce dataset. Passages aligning with the reference answer are highlighted in the same color. The left side of the figure lists the scores for each baseline.
Appendix F Prompts Template
We provide the prompt templates for all multi-agent baselines presented in the paper, exemplified using the automotive business workflow.
F.1 Cochain
We present a specific overview of the prompts template in the Cochain. As illustrated in Figure 14, the prompt template adopts a modular design comprising three core components: the knowledge module retrieved from the collaborative knowledge graph, the causal chain module, and the business workflow prompts module generated through prompts tree retrieval. In the template, the gray box represents fixed prompt content, while the colored boxes denote dynamic content that adapts based on user needs.
Figure 14: The prompts template of Cochain for final input to LLM.
F.2 PMC
Figure 15: The manager agent prompts template of PMC.
Figure 16: The supervisor agent prompts template of PMC.
Figure 17: The deliverer agent prompts template of PMC.
Figure 18: The executor agent prompts template of PMC.
F.3 MedAgents
Figure 19: The prompt templates for stages 1 through 3 of MedAgents.
Figure 20: The prompt templates for stages 4 and 5 of MedAgents.
F.4 Debate
Figure 21: The prompts template of Debate for final input to LLM.
F.5 CoA
Figure 22: The prompt templates for the worker agent and manager agent of CoA.
Figure 23: The prompt templates for each worker agent on the Automotive dataset.